### v1

# Benchmarking the Spatial Robustness of DNNs via Natural and Adversarial Localized Corruptions

Abstract
The robustness of DNNs is a crucial factor in safety-critical applications, particularly in complex and dynamic environments where localized corruptions can arise. While previous studies have evaluated the robustness of semantic segmentation (SS) models under whole-image natural or adversarial corruptions, a comprehensive investigation into the spatial robustness of dense vision models under localized corruptions remained underexplored. This paper fills this gap by introducing specialized metrics for benchmarking the spatial robustness of segmentation models, alongside with an evaluation framework to assess the impact of localized corruptions. Furthermore, we uncover the inherent complexity of characterizing worst-case robustness using a single localized adversarial perturbation. To address this, we propose region-aware multi-attack adversarial analysis, a method that enables a deeper understanding of model robustness against adversarial perturbations applied to specific regions. The proposed metrics and analysis were evaluated on 15 segmentation models in driving scenarios, uncovering key insights into the effects of localized corruption in both natural and adversarial forms. The results reveal that models respond to these two types of threats differently; for instance, transformer-based segmentation models demonstrate notable robustness to localized natural corruptions but are highly vulnerable to adversarial ones and vice-versa for CNN-based models. Consequently, we also address the challenge of balancing robustness to both natural and adversarial localized corruptions by means of ensemble models, thereby achieving a broader threat coverage and improved reliability for dense vision tasks.
1 Introduction
In recent years, deep neural networks (DNNs) have demonstrated remarkable performance across various vision applications, increasing their potential applicability in safety-critical domains such as autonomous systems[19, 27]. Among these scenarios, dense prediction tasks like semantic segmentation (SS) have garnered significant attention. These tasks require models to understand the semantic meaning of each pixel in a given scene, often necessitating the extraction of deep contextual information. In this context, it is crucial to assess the robustness and trustworthiness of SS models, as complex and dynamic environments can give rise to unexpected input variations that challenge the entire system reliability.
Despite this importance, the complexity of DNNs and real-world application scenarios, as in autonomous and assisted driving, has made ensuring and evaluating robustness one of the most challenging problems in modern AI [6, 3, 32, 16]. In the literature, robustness evaluation typically involves analyzing model performance under corrupted input conditions, with two main types of corruption commonly studied: (i) Natural corruptions, which simulate faults or real-world environmental conditions such as sensor noise or adverse weather (e.g., snow, rain) [14]; (ii) Adversarial corruptions, which involve intentionally crafted perturbations designed to manipulate model predictions. While adversarial attacks are often studied as a security threat, they are also a valuable tool for evaluating a model’s robustness under a close approximation of worst-case scenarios, contributing to a sense of verifiability [4, 16].
In recent years, numerous studies have evaluated the robustness of vision models against both natural and adversarial perturbations [14, 32, 17]. However, a comprehensive characterization of their behavior under localized corruptions [18, 29, 30] remains insufficiently explored, particularly for dense prediction tasks such as semantic segmentation. In fact, although poorly explored, localized perturbations are highly relevant to real-world scenarios (e.g., sunlight affecting specific regions of an input image) and need to gather more attention in robustness studies. This gap raises critical questions about the spatial robustness of segmentation models: To what extent do corrupted regions influence predictions in both the corrupted and uncorrupted areas of an image? How effectively can models leverage uncorrupted regions to mitigate the impact of corrupted ones? Addressing these questions is crucial for advancing the robustness and reliability of scene-understanding models in safety-critical environments.
This paper addresses the above challenges by studying the effect of localized corruptions across different models through the lens of both natural transformations and adversarial perturbations. In particular, the work begins by formalizing the concept of localized corruptions and proposing ad-hoc metrics specifically designed to evaluate spatial robustness in different regions of the input for SS models. We then introduce an analysis framework for benchmarking the spatial robustness of SS models against localized natural corruptions, enhancing the understanding of the reliability of DNNs in practical scenarios, where localized corruptions can occur, as in driving environments.
Next, we investigate critical aspects of adversarial attacks within the context of SS tasks, highlighting that identifying a single localized perturbation capable of comprehensively evaluating spatial robustness (e.g., misclassifying as many pixel predictions as possible) is particularly challenging due to the multi-objective nature of the adversarial optimization problem. To address this challenge, we propose region-aware multi-attack adversarial analysis, a method that iteratively generates localized perturbations, each targeting regions previously unaffected by prior attacks. The purpose of this method is to demonstrate that regions initially considered robust (as they are not misclassified by a single localized attack) can indeed be compromised when specific attack configurations are applied. This insight enhances the understanding of worst-case localized perturbations and their impact on model robustness, highlighting the need to understand the extent to which pixels can be perturbed when a perturbation is applied to a specific area.
The proposed framework, considering both localized natural and adversarial corruptions, is then applied in an experimental evaluation to benchmark the spatial robustness of $15$ SS models in driving scenarios [8]. Most interestingly, we observed that lightweight convolution-based SS models are significantly vulnerable to localized natural corruptions, while Transformer-based models exhibit higher robustness. Conversely, the former are more robust to adversarial perturbations, while the latter and highly vulnerable to the same. These findings raise significant concerns about the reliability of SS models when evaluated under worst-case safety-related scenarios.
To address this issue and highlight the importance of achieving a trade-off in terms of spatial robustness between natural transformations and adversarial perturbations, we finally explore the design and configuration of ensemble strategies.
Overall, this study marks a significant step toward understanding localized corruptions and paves the way for developing more robust SS models in practice.
Contribution. In summary, this paper makes the following contribution:
  • We formalize new metrics to evaluate the accuracy of SS models while accounting for their spatial robustness under localized corruptions. Furthermore, we extend existing evaluation approaches to study localized corruption by proposing a flexible framework that allows for quick evaluation of the model spatial robustness.
  • We uncover important insights into the study of localized adversarial perturbations for spatial evaluation in SS models and introduce a new method for generating adversarial perturbations named region-aware multi-attack adversarial analysis.
  • We applied the proposed metrics, framework, and method to analyze the spatial robustness of 15 SS models under natural and adversarial localized perturbations, revealing contrasting behaviors between the two types of corruptions. Finally, we study the trade-off between robustness to natural and adversarial corruptions by means of ensemble models.
The rest of the paper is structured as follows: Section 2 discusses related work and remarks the importance of studying the spatial robustness of dense prediction models; Section 3 introduces the proposed analysis and formalizes the metrics we used; Section 4 defines the evaluation framework adopted for assessing models under localized natural corruptions; Section 5 focuses on localized adversarial perturbations; Section 6 discusses our experimental evaluations and ensemble models, and finally, Section 7 states the conclusions and discussion.
2 Related Work
Analysis and benchmarks on DNNs robustness
Deep learning models are widely used across various applications, including safety-critical domains such as autonomous driving. Given their increasing deployment in such scenarios, assessing their robustness is essential to evaluate their reliability and suitability for real-world use. To address these challenges, several benchmarks and robustness analyses have been proposed over the years to assess model reliability against both natural and adversarial corruptions. Notable examples include [14, 10, 17, 31].
Regarding natural corruptions, the objective is to evaluate models’ robustness against perturbations that simulate adverse or potentially out-of-distribution conditions that can be encountered in real-world environments, such as noise, rain, or snow. One of the earliest and most influential studies in this area was conducted by Hendrycks and Dietterich, who introduced a benchmark to assess the resilience of image classification models. This benchmark comprises various natural corruptions, including weather-related disturbances and image formatting artifacts [14]. Since then, these corruptions have been widely adopted and extended in subsequent studies to systematically evaluate and compare the robustness of different models [13, 2, 25]. For instance, Bhojanapalli et al. [2] investigated the robustness of ViTs comparing it to the robustness of a ResNet-50 model under image corruptions, concluding that ViTs models are at least as robust as the CNN take into consideration on the majority of the corruptions. While the majority of studies have focused on classification tasks, fewer works have explored the robustness of models in dense prediction tasks, such as semantic segmentation [31, 17], which have highlighted the need for more comprehensive evaluations in this new domain
Beyond natural corruptions, other works have examined adversarial corruptions [32, 3], which represent a worst-case scenario from a security perspective, where perturbations are deliberately crafted to deceive models. Similar to natural corruptions, several studies have compared the robustness of different architectures against adversarial perturbations [2, 25, 22]. Specifically, research on adversarial robustness in semantic segmentation [37, 1, 17] has underscored the importance of establishing more rigorous benchmarks and developing more effective defense strategies to enhance model resilience in real-world applications.
Evaluation of localized corruptions and spatial robustness
While most of the aforementioned studies primarily focus on corruptions affecting the entire input, it is crucial to consider the problem of spatial robustness [30, 29, 13]. This aspect is particularly relevant for dense prediction tasks, where model predictions may be influenced by regions beyond the immediate locality of the target area. For instance, in semantic segmentation, pixels that are spatially distant from the region of interest can still impact the model’s decision in unexpected ways and vice versa.
Despite its importance, research on spatial robustness under natural corruptions remains limited. For example, [13] analyzes the robustness of ViTs compared to CNNs under natural and adversarial corruptions concluding that ViTs generally exhibit greater resilience to natural corruptions than CNNs. However, does not specifically address dense prediction tasks. Conversely, most existing studies on spatial robustness have focused on adversarial settings, particularly adversarial patches [26, 5, 20, 23], while only a few works have explored the effects of localized adversarial perturbations [24]. Additionally, there is still a lack of comparative analysis on transformer-based models in both cases.
In this study, we primarily focus on localized adversarial perturbations as they are more directly related to worst-case safety evaluations. While adversarial patches pose a significant security concern, their effects and mitigation strategies have been explored in the literature [28, 13]. Our work aims to bridge this gap by also investigating the unexplored case of natural corruptions, which is particularly relevant for safety assessments. We provide a more comprehensive analysis of spatial robustness in the context of adversarial perturbations, with a specific focus on comparing transformer-based and convolution-based architectures.
3 Spatial Robustness Analysis
This section first provides some preliminary and background information to theoretically define localized perturbation within the context of SS models. Then, new metrics for spatial robustness are proposed to generalize the classic pixel-wise accuracy score used in semantic segmentation.
3.1 Background and Formalisms
Preliminaries
We consider input images of size $H\times W$ and denote by $\mathcal{I}$ the set of their pixels, which are then individually referred to using the index $i$. A semantic segmentation model designed to classify each pixel into one of $N$ classes is represented by a function $f:x\mapsto\mathbb{R}^{(H\cdot W)\times N}$, which outputs a per-class distribution of scores for each pixel $i$. The vector of output scores for each pixel $i$ is denoted by $f_{i}$ while its component related to the $j$-th class is denoted by $f^{j}_{i}$. The semantic label predicted for pixel $i$ is determined by selecting the class with the highest score for that pixel, i.e., $\operatorname*{argmax}_{j\in\{1,\dots,N\}}f_{i}^{j}(x).$ Finally, the corresponding ground-truth segmentation map for the image $x$ is denoted by $y\in\mathbb{I}^{H\times W}$, where each pixel is assigned its true class label, i.e., $y_{i}\in\{1,\dots,N\}$.
Corruption Areas and Ratio
To test the spatial robustness of SS models, we consider corruptions applied to only part of the image. Specifically, following the notation adopted in previous work [24, 29], we define a corrupted area $M$ as a binary mask in the input space $M\in\{0,1\}^{H\times W}$, which indicates where the corruption is applied. For any 2D pixel with index $i$, $M_{i}=1$ denotes the pixel is corrupted, while $M_{i}=0$ denotes that it is not. The corrupted image $x_{M}^{c}$, given a mask $M$ and a corruption function $c(\cdot)$, is computed as:
$x^{c}_{M}=c(x)\cdot M+x\cdot(1-M),$ (1)
where $c(\cdot)$ can represent either a natural transformation or an adversarial attack. Additionally, we introduce the ratio of the perturbed area $r=\sum_{i}M_{i}/|M|$ and the non-corrupted region mask, denoted by $\overline{M}$, which is the complement of $M$, i.e., $\overline{M}=1-M$.
3.2 Metrics for Spatial Robustness
Next, we define metrics for studying the effect of localized corruptions on predictions in different areas of the image, by generalizing both the classic accuracy metrics used for SS and those applied for testing the robustness of image-classification DNNs against corruptions.
Classic Corruption Accuracy
Deriving from the analysis of [14] in image classification, given a model $f$ and a set of possible corruptions $\mathcal{C}$ (e.g., Gaussian noise, adversarial perturbation or None to denote no corruptions), the robustness accuracy of an SS model on a dataset $\mathcal{D}$ is defined as:
$A(f,\mathcal{C})=\frac{1}{|\mathcal{C}|\cdot|\mathcal{I}|\cdot|\mathcal{D}|}\sum_{c\in\mathcal{C}}\sum_{(x,y)\in\mathcal{D}}\sum_{i\in\mathcal{I}}\mathds{1}\big{(}f_{i}(c(x))=y_{i}\big{)},$ (2)
where $|\mathcal{C}|$ is the number of corruption types, $|\mathcal{I}|$ is the number of pixels in the input, $|\mathcal{D}|$ is the size of the dataset, and $\mathds{1}(\cdot)$ is an operator that returns ’1’ if a condition is ’True’, otherwise ’0’. Following this metric, we can also define the Corruption Error as:
$\textit{CE}(f,\mathcal{C})=\frac{A_{\textit{base}}-A(f,\mathcal{C})}{A_{\textit{base}}},$ (3)
where $A_{\textit{base}}$ refers to the accuracy of a reference model (either $f$ or another one, e.g., AlexNet, as used in [14]). This enables a better understanding of the impact of a given corruption with respect to a reference value.
Localized Corruption Accuracy
We extend the analysis of the accuracy to consider the case of localized corruption, identified by a corruption mask $M$. The accuracy of a model $f$ under a corruption $c$ applied in the areas denoted by $M$ is defined as:
$A_{\mathcal{P}}(f,\mathcal{C},M)=\frac{1}{|\mathcal{C}|\cdot|\mathcal{D}|}\sum_{c\in\mathcal{C}}\sum_{(x,y)\in\mathcal{D}}\mathcal{K}_{\mathcal{P}}(f,x,c,M,y),$ (4)
with
$\mathcal{K}_{\mathcal{P}}(f,x,c,M,y)=\sum_{i\in\mathcal{I}}\mathds{1}\big{(}f_{i}(c(x))=y_{i}\big{)}\cdot\mathcal{P}_{i}(M),$ (5)
where $\mathcal{P}_{i}(M)$ is a spatial importance function designed to weigh the significance of misclassifying a pixel $i$ based on the corruption mask $M$. This function plays a critical role in assessing the impact of regions of the input image, allowing the analysis to either focus on regions of higher relevance or those affected by corruption. While, in general, spatial importance functions can be arbitrary defined, in this study we focus on two specific definitions:
  • Non-Corrupted Region Analysis ($A_{\overline{M}}$): this approach focuses on evaluating the spatial robustness of the model by analyzing regions unaffected by corruptions, uniformly weighting all non-corrupted pixels, i.e. $\mathcal{P}_{i}(M)=\overline{M}_{i}/\left|\overline{M}\right|$.
  • Corrupted Region Analysis ($A_{M}$): this approach is the dual of the one above, focusing on corrupted pixels only, i.e., $\mathcal{P}_{i}(M)=M_{i}/|M|$.
The first definition ($A_{\overline{M}}$) is particularly useful for assessing how well the model remains accurate in non-corrupted regions, offering insights into the capability of corruptions to extend beyond the actual corrupted region. Conversely, the second definition ($A_{M}$) focuses on the model’s ability to recover accurate predictions in corrupted regions, potentially leveraging information from non-corrupted regions ($\overline{M}$). This analysis is more meaningful when studying natural corruptions, as adversarial attacks can easily lead to complete misclassification of corrupted regions [29, 24], making robustness assessments in such areas less informative. In our experimental analysis, we highlight that these metrics are expressive for both natural and adversarial corruptions, as they provide measures of resiliency to corruptions within and far from the targeted regions.
It is worth noting that the metrics presented above can be computed for multiple corruption regions: different settings are considered in our experimental evaluation (Section 6). For localized adversarial perturbations, we follow traditional approaches that focus on statically targeting specific areas of the image, such as the center or a corner (e.g.,  [29, 24]). The rationale behind this choice is that these positions allow evaluating the worst- or best-case scenarios of a model’s receptive field [21].
Different considerations can be made when assessing the impact of arbitrarily-placed localized corruptions. For natural corruptions, we consider the above metrics while selecting the corruption regions randomly, leading to the following formulation:
$A_{\mathcal{P}}(f,\mathcal{C})=\frac{1}{|\mathcal{C}|\cdot|\mathcal{D}|}\sum_{c\in\mathcal{C}}\sum_{(x,y)\in\mathcal{D}}\mathcal{K}_{\mathcal{P}}(f,x,c,M_{r}(x),y),$ (6)
where $M_{r}(x)$ is a mask that determines the corrupted regions, with patches randomly selected for each input $x$ while respecting a ratio $r=\sum_{i}M_{i}/|M|$ of the corrupted area.
It is finally important to remark that all the proposed metrics focus on pixel-wise accuracy for semantic segmentation. Other metrics, such as the mean intersection over union (MIoU), are also frequently used in the field of SS and our analysis can be extended to support them. For instance, no significant modifications are required to support MIoU.
4 Localized Natural Corruptions
Figure 1: Illustrations of the proposed framework for evaluating localized natural corruptions. A configuration is specified to define the settings for the localized corruption analysis, and a single sample from the dataset is returned as a tuple consisting of the original image, the original ground truth label, and the corrupted image, along with the corrupted region that highlights which areas have been perturbed.
This section presents the framework we designed to evaluate localized natural corruptions. Our framework was inspired by ImageNet-C [14], which however does not support localized corruptions. We consider different transformations $c$, which can be applied at various severity levels $s$ (i.e., 1, 2, 3, 4, 5), as selected by the user. A natural corruption applied to an input $x$ with severity level $s$ is denoted by $c(x)=\Gamma(x,s)$, where $\Gamma$ represents a natural transformation, such as Gaussian noise, blur, motion blur, etc.
The corrupted region $M$ is automatically selected according to the following pipeline. The image $x$ is first divided into non-overlapping patches with size $\Delta=(\Delta_{x},\Delta_{y})$, thereby obtaining a total of $P=P_{x}\times P_{y}$ patches with $P_{x}=\lceil W/\Delta_{x}\rceil$, $P_{y}=\lceil H/\Delta_{y}\rceil$, where $H$ and $W$ are the height and width of $x$, respectively. Given a specified corruption ratio $r$, each patch, and therefore all its pixels, is selected to be perturbed with probability $r$. All pixels within a selected patch are assigned a value of $1$ in $M$, while the others have a value of 00.
In Figure 1, we present an illustration of a possible configuration used within the tested framework and the output of a dataset sample in the corrupted validation set of Cityscapes [8]. Specifically, a sample corresponds to a tuple containing the original image and its label (kept available to allow comparisons during the evaluation), along with the perturbed image and the corruption region mask $M$. During evaluation, for each image, the set of patches to be perturbed (and thus the mask $M$) is recomputed, effectively randomizing the selection of the perturbed areas for the evaluation.
5 Localized Adversarial Corruptions
This section presents our novel analysis for localized adversarial perturbations. Although previous work already addressed the study of localized adversarial perturbations against SS models [24], we next highlight the difficulties of performing a comprehensive spatial adversarial evaluation using a single attack. To this end, we begin by discussing classic approaches and pointing out their drawbacks. Finally, we introduce our region-aware multi-attack algorithm designed to address these issues and offer a better understanding of the worst-case spatial robustness of ML models.
5.1 Attacks for SS from previous work
In the context of classic adversarial perturbations [32, 3], the main idea is to craft a specific adversarial noise $\delta$ that realizes a worst-case perturbation constrained by a magnitude $\varepsilon$ under a specific norm, e.g., the $l_{\infty}$ norm, such that $\|\delta\|_{\infty}\leq\varepsilon$. Adversarial perturbations are typically computed by optimizing an adversarial loss function $\mathcal{L}$, e.g., cross-entropy in image classification, which aims to find a perturbation $\delta$ for an input $x$ that forces the model to make incorrect predictions $\neq y$:
$\delta=\arg\max_{\|\delta\|_{\infty}\leq\varepsilon}\mathcal{L}(f(x+\delta),y),$ (7)
After recalling Equation 1, it is possible to express inputs affected by localized attacks with a perturbation $\delta$ and a mask $M$ as follows:
$x_{M}^{\delta}=(x+\delta)\cdot M+x\cdot(1-M).$ (8)
When interested in dense prediction tasks, such as semantic segmentation, where predictions are made at the pixel level, the perturbation is generally computed to maximize the misclassification of all or a subset of pixels in the image [24, 29]. In the latter case, it is possible to define a fooling region as a subset of the input space for which we intend to induce a misprediction. The fooling region can be arbitrarily defined by a mask $F\in\{0,1\}^{H\times W}$ that does not necessarily need to overlap with the corruption mask $M$. In this context, the perturbation $\delta$ can be optimized as follows:
$\delta=\arg\max_{\|\delta\|_{\infty}\leq\varepsilon}\mathcal{L}_{\mathcal{F}}(f(x^{\delta}_{M}),y),$ (9)
where $\mathcal{L}_{\mathcal{F}}$ is an adversarial loss designed to focus the model performance in the fooling region $F$ only. In practice, to solve the above optimization problem, the perturbation $\delta$ can be iteratively updated with step size $\alpha$ to maximize the classification error:
$\delta_{i+1}\leftarrow\delta_{i}+\alpha\cdot\text{sign}\left(\nabla_{\delta}\mathcal{L}_{\mathcal{F}}\right).$ (10)
5.2 Limitations of previous work
Although previous approaches are capable of crafting strong perturbations by corrupting as many pixels as possible, we argue that comprehensively evaluating the spatial robustness of semantic segmentation models requires multiple instances of attacks applied to the same area. A single perturbation, even when computed using a strong attack, cannot cover all regions susceptible to misprediction. This happens because, as shown in Eq. 9, crafting an optimal localized adversarial perturbation involves solving a complex multi-objective optimization problem, where the goal is to increase the loss for each pixel in the image, representing a wide range of distinct objectives. Many of these objectives are difficult to solve simultaneously, often leading to suboptimal or locally optimal solutions. To support this observation, we provide representative analysis results in Figures 2(a) and  2(b) (discussed in the next paragraphs), which highlight, from both inter-class and intra-class perspectives, that adversarial attacks targeting specific subsets of pixels may not generalize effectively to others, making it particularly challenging to craft a single localized perturbation that generates worst-case adversarial effects.
(a) Inter-class target attack analysis.
(b) Intra-class target attack analysis
Figure 2: Analysis of limitations in SS adversarial optimization for single localized perturbations: (a) Inter-class analysis: Attacked images and predictions across different target classes. (b) Intra-class analysis: Targeted multi-attack on the class ‘Car’. Visualization of adversarial effects in different areas. Each subfigure highlights a distinct aspect of the attack.
Let us first focus on inter-class attacks with reference to Figure 2(a), where localized perturbations are applied to the center of the image ($100\times 100$, using iterative targeted attacks [24] with $\epsilon=32/255$). In this example, attacks are performed each targeting different classes in separate runs (’road’, ’sky’, and ’car’). The outputs of the addressed targeted attacks and their corresponding classes are shown from the third to fifth subfigure. Additionally, we also show the output prediction of an untargeted attack (targeting all classes), in the last subfigure. The results demonstrate that the adversarial patterns produce distinct adversarial effects, highlighting that the multi-objective optimization landscape for different target classes is highly divergent, evidencing difficulties for a targeted attack to affect the areas interested by others. Furthermore, when observing the output of the untargeted attack, it closely resembles the output of the ’road’ attack, suggesting a bias in the multi-objective optimization process, possibly given by the fact that the majority of pixels of the image are classified as ’road’.
We then analyze the same problem in the context of intra-class attacks. In this case, as shown in Figure 2(b), we crafted attacks to misclassify pixels belonging to the ’car’ class. We started with an initial fooling region, highlted in the first subfigure, consisting of all pixels labeled as ’car’ in the ground truth. After executing the attack of Eq. (9) a first time (denoted by attack 1) using the initial fooling region, we were capable of misclassifying a limited subset of the targeted pixels only (note the attack output in the second row and the corresponding red pixels in the third row of Figure 2(b)). To carry out a second attack (attack 2), we excluded from the initial fooling area the pixels that were successfully attacked in the previous attack. This adjustment allows us to investigate whether previously attacked pixels negatively influence the optimization process for the remaining pixels classified as ’car’. The restricted fooling region for the second attack is shown in the third row. As it can be observed from Figure 2(b), the second attack expands the attacked area, indicating that previously attacked regions can indeed hinder the optimization process when exploring new attackable regions. This process is repeated for a third attack (last columns in the second and third rows of the figure), confirming the previous observation.
Finally, for completeness, we present the cumulative adversarial output in the third subfigure (first row, third column). This cumulative output assigns each pixel its adversarial misclassified prediction if at least one attempted attack has misclassified it (the first attack is considered for simplicity); otherwise, it retains the original model prediction. This approach provides a clearer worst-case understanding of whether each given pixel can be misclassified by at least one attack and will be used in the algorithm proposed in the next subsection.
5.3 Achieving Stronger Spatial Adversarial Analysis
The observations made in the previous section allow concluding that single instances of SS attacks provide an incomplete assessment of the worst-case spatial adversarial robustness of a model. To address this limitation and establish a more comprehensive benchmark for evaluating SS models under localized adversarial perturbations, we propose a new method, called region-aware multi-attack, in Algorithm 1. This algorithm builds upon the observations made above and is next explained step-by-step. At a high level, the method iteratively attacks the input while progressively modifying the attacked area by refining the fooling region at each attack run, thereby maximizing the coverage of attacked pixels.
Algorithm 1 Region-aware Multi-Attack Analysis
1:
2:Model $f$; input $x$; labels y; Num. of attacks $N_{\text{Att}}$
3:$\hat{y}_{\text{clean}}\leftarrow\text{argmax}(f(x))$triangleright Clean preds
4:$\hat{y}_{\text{O}}\leftarrow\hat{y}_{\text{clean}}$triangleright Initialize cumulative output
5:$M_{\text{corr}}\leftarrow(\hat{y}_{\text{O}}=\textbf{y})$triangleright Mask of correct preds
6:$\mathcal{F}\leftarrow M_{\text{corr}}$triangleright Initial fooling region
7:for $i\leftarrow 1$ to $N_{\text{Att}}$ do
8:     $\mathcal{F}\leftarrow\mathcal{F}\cdot M_{\text{corr}}$triangleright Update fooling region
9:     $x_{\text{adv}}\leftarrow\textit{Attack}(x,\textbf{y},f,\mathcal{F})$
10:     $\hat{y}_{\text{adv}}\leftarrow\text{argmax}(f(x_{\text{adv}}))$triangleright Preds after attack
11:     $M_{\text{cur}}\leftarrow(\hat{y}_{\text{adv}}=\textbf{y})$triangleright Current correct pixels
12:     $\bar{M}\leftarrow(1-M_{\text{cur}})\cdot M_{\text{corr}}$triangleright Newly misclassified pixels
13:     $M_{\text{corr}}\leftarrow M_{\text{cur}}\cdot M_{\text{corr}}$triangleright Update correct preds mask
14:     $\hat{y}_{\text{O}}\leftarrow M_{\text{corr}}\cdot\hat{y}_{\text{O}}+\bar{M}\cdot\hat{y}_{\text{adv}}$triangleright Update cumul. output
15:end for
16:return $\textit{Eval}(\hat{y}_{\text{O}},\textbf{y})$
The proposed algorithm takes as input a given image $x$, its labels y, and an SS model $f$, and aims to evaluate the robustness of the image by running multiple instances of attacks ($N_{\textit{Att}}$) based on a given attack method (referred to as Attack in the algorithm). At the beginning of the algorithm, the cumulative output is initialized to the correct model predictions, and the initial fooling region consists of all pixels that are correctly predicted in the non-perturbed output (line 4) [^1] We address the untargeted case for simplicity and to generalize as much as possible, covering as many pixels as possible. However, the approach can be easily extended to a targeted formulation.. Subsequently, the attack iteratively targets the pixels in the fooling region using a classic iterative attack (e.g., Equation 10, line 7), computes the corresponding output (line 8), and updates the fooling region by identifying pixels in the preceding fooling region that have not yet been misclassified in the current attack (lines 9–12 and line 6). At the end of each attack, the cumulative adversarial output is updated to replace the predictions of newly misclassified pixels that had previously been correctly classified in earlier attacks.
Finally, the cumulative output is evaluated according to the desired metrics (Eval in the algorithm). It is important to remark that the cumulative output is not derived from the model’s prediction on a single perturbed sample but rather from the cumulative worst-case attacked outputs, obtained by aggregating misclassifications across multiple localized perturbations, all concentrated in the same area. From a safety perspective, this approach provides a stronger and more accurate understanding of the worst-case robustness of each pixel, demonstrating the existence of at least one perturbation capable of misclassifying them.
Please note that the proposed formulation is entirely untargeted. As demonstrated in the experimental section, the approach also addresses class-wise coverage by incorporating an incremental fooling region. This strategy helps mitigate intra-class issues that may arise even in targeted attacks, as discussed above and illustrated in Figure 2(b).
6 Evaluation of Spatial Robustness
In the following, we first describe the experimental setup. Then, we evaluate the spatial robustness, based on the proposed metrics, of several models against localized natural corruptions and localized adversarial attacks separately. Finally, we discuss the importance of addressing these two aspects together and propose an ensemble-based analysis aimed at balancing localized natural and adversarial robustness.
6.1 Experimental Setup
The experiments were conducted with NVIDIA A100 GPUs, using the validation set of Cityscapes [9], which serves as the reference driving dataset for high-resolution scene-understanding segmentation. In the context of semantic segmentation, this dataset has frequently been used for robustness evaluation, as it addresses real-world and complex outdoor scenarios [29, 24, 28]. The Cityscapes validation set includes 500 high-resolution images with an original resolution of ($1024\times 2048$), resized to manage computational costs (as done in [24]). We used this set to evaluate both localized natural corruptions and adversarial attacks.
Models. We targeted multiple models for semantic segmentation, all known for their ability to achieve high performance on large images while maintaining practical inference times affordable for many real-time applications. Specifically, we considered ICNet [39], BiseNet [36], DDRNet [15], PSPNet [40], SegFormer [34], PIDNet [35], and DeepLabV3 [7], including multiple versions of each model to evaluate robustness differences, such as variations in the backbone architecture. The average accuracy of these models on the clean validation set of Cityscapes is reported in the first column of Table 1.
Metrics. The spatial robustness analysis was conducted by evaluating the pixel-wise accuracy of the models in the con-corrupted region $A_{M}$ and the corrupted region $A_{\overline{M}}$, both defined in Section 3. Furthermore, to better understand the robustness degradation as the evaluation parameters vary, we analyzed the previous scores in terms of the relative corruption error (RCE) with respect to the accuracy of the addressed models in the same area of interest when no natural corruption or attacks are applied. Leveraging the localized corruption accuracy defined in Eq. (4), in the case of the relative corruption error in the corrupted region the relative error can be computed as
$\text{RCE}(f,C,M)=\frac{A_{\mathcal{P}}(f,\{\emptyset\},M)-A_{\mathcal{P}}(f,C,M)}{A_{\mathcal{P}}(f,\{\emptyset\},M)},$
were $\mathcal{P}_{i}(M)=M_{i}/|M|$. A similar definition can be derived for the RCE outside the corrupted region. Note also that, as indicated in Section 3, natural corruptions have been evaluated across different regions (see Equation (6)).
Type Brightness Snow Gaussian Noise
Severity Clean 1 3 5 1 3 5 1 3 5
Metric $M$ $\overline{M}$ $M$ $\overline{M}$ $M$ $\overline{M}$ $M$ $\overline{M}$ $M$ $\overline{M}$ $M$ $\overline{M}$ $M$ $\overline{M}$ $M$ $\overline{M}$ $M$ $\overline{M}$
BisenetX39 70.6 63.1 67.0 47.0 61.4 34.7 54.9 34.9 57.4 22.1 51.4 20.6 51.3 50.0 61.5 20.4 50.9 12.1 43.4
BisenetR18 78.0 72.6 76.0 59.5 71.5 49.1 67.9 42.2 68.7 24.7 64.2 19.7 62.0 66.4 73.9 42.4 67.6 21.5 62.4
BisenetR101 82.3 76.8 80.3 63.7 77.3 53.7 74.1 39.7 72.5 20.8 68.2 19.8 68.2 64.0 78.5 35.8 74.3 22.4 70.7
DDRnet23 83.2 78.7 81.7 64.7 76.4 51.2 72.5 39.2 71.8 22.5 67.4 18.5 65.1 64.2 78.1 27.3 67.9 13.7 59.8
DDRnet23Slim 80.8 75.0 78.9 61.4 73.0 49.4 68.8 39.9 68.2 25.4 63.5 20.1 60.4 63.5 74.1 32.9 64.4 15.0 58.3
Icnet 72.7 65.8 70.4 50.3 64.0 41.4 58.6 40.9 58.7 22.7 48.1 20.4 47.6 59.2 67.3 37.5 58.1 20.8 47.8
PSPnet 81.6 75.1 79.1 60.5 74.5 49.1 71.6 35.2 70.1 14.6 63.6 11.6 60.4 66.8 78.2 35.3 70.9 15.6 59.8
Segformer-bo 79.3 75.8 78.2 67.6 76.4 62.0 75.7 45.9 73.2 26.5 70.9 19.6 69.9 68.8 77.1 54.0 74.8 34.8 73.0
Segformer-b1 80.9 77.4 80.2 70.2 79.1 66.0 78.6 50.1 77.3 27.9 75.4 21.1 74.5 72.5 80.2 60.2 79.1 42.2 77.7
Deeplabv3-m 76.7 73.3 75.0 66.6 72.3 59.2 69.4 50.5 69.8 32.1 63.9 25.2 59.9 56.4 71.4 36.3 64.4 20.2 57.6
Deeplabv3-r 80.6 77.0 78.6 71.7 75.7 64.8 73.0 49.3 72.6 29.2 68.4 22.3 66.1 65.8 76.4 40.3 67.6 16.3 58.1
Pidnet-s 80.0 74.1 77.1 61.7 73.4 48.8 69.0 41.7 69.5 22.3 63.9 18.2 61.2 65.6 76.3 34.3 66.2 12.2 52.8
Pidnet-m 80.6 75.9 79.2 62.9 75.2 48.9 71.2 47.6 73.5 28.1 69.6 22.6 67.1 59.4 75.2 27.3 68.8 10.1 64.4
Pidnet-l 82.1 77.2 81.1 60.9 76.7 47.3 72.4 47.7 74.3 24.9 69.3 19.0 67.4 69.0 79.5 36.4 71.7 15.6 63.3
Table 1: Analysis of natural corruptions considering brightness, snow, and gaussian noise as corruption types at different severity levels (1, 2, 3). For comparison, the first row reports the clean average accuracy across the entire image (MeanAcc score), while the remaining columns analyze the accuracy within and outside the corrupted region, represented as $M$ and $\overline{M}$, respectively, following the analysis presented in Section 3. The corruption ratio is fixed at 50% of the image, and the same seed is applied for each test to ensure fair comparisons. For each metric and analysis, the top-2 models are highlighted in green for easier interpretation.
Figure 3: Illustrations of the effects of localized natural corruptions across different semantic segmentation models (BiseNetX39, PSPNet, SegFormer1, DeepLabV3-ResNet) for localized natural corruptions with a ratio of $30\%$ and severity level 3. Synthetic snow is shown in the third row, and Gaussian noise in the second row, while the clean image and corresponding outputs are displayed in the first row.
6.2 Evaluation of Natural Robustness
We first evaluate the natural robustness of the selected SS models. Table 1 reports the results for the Non-Corrupted Region Analysis ($\overline{M}$ in the table) and the Corrupted Region Analysis ($M$ in the table), considering different localized natural corruptions (synthetic snow, brightness adjustments, and gaussian noise) across different severity levels. For these tests, the corruption ratio is fixed to $r=0.5$, and corruptions are applied by splitting the images into patches of size $(256,256)$. To ensure fairness in the evaluation process, the seed was consistently initialized for each test. To improve table readability, we highlighted the top-2 models for each column.
Considering the robustness in the corrupted region (columns labeled with $M$), for snow and brightness adjustments, no model demonstrates a clear and consistent superiority over the others, suggesting that convolution-based architectures remain valid alternatives to SegFormers. This observation does not hold in the case of gaussian noise, where transformer-based architectures exhibit a significant accuracy advantage in the corrupted region. Regarding the robustness in the non-corrupted region (columns labeled with $\overline{M}$), transformer-like architectures achieve significnatly better performance, especially when addressing higher severity levels of localized corruptions. In these cases, natural corruptions have the potential to substantially affect non-corrupted regions, thereby reducing the accuracy of $A_{\overline{M}}$. However, SegFormer demonstrates robustness by preventing the spread of mispredictions to otherwise clean regions.
In Figure 3, we also show the outputs of the models when localized natural perturbations are applied, while the clean image and corresponding outputs of different models are displayed in the first row. For this test, we use a corruption ratio of $r=0.3$ and a severity level of $3$. The corresponding results align with those in the table and previous observations, showing that, in general, the SegFormer architecture achieves better robustness both within and outside the corrupted region.
(a) Synthetic snow analysis
(b) Gaussian noise analysis
Figure 4: Analysis of the impact of the corruption ratio. We report the relative error with respect to the non-corrupted case, evaluated within the corrupted region (left plots) and the non-corrupted region (right plots). In the top plots, we use synthetic snow with severity level 3, and in the bottom plots, we use Gaussian noise with the same severity level.
Impact of the corruption ratio
To further analyze the impact of localized corruption, we report in Figure 4(a) and Figure 4(b) the variation of the RCE (computed with respect to the clean accuracy in the addressed region, as described in the experimental setup) as a function of the corruption ratio. The RCE was computed both within (left plots) and outside (right plots) the corrupted regions, considering synthetic snow and Gaussian noise, respectively.
As shown in the plots, the corruption error follows different trends across the models. For small corruption ratios and corruption types with limited impact, such as synthetic snow, convolution-based models such as DeepLab-MobileNet and DeepLab-ResNet achieve performance comparable to SegFormer, both within and outside the corrupted region. However, when the corruption ratio increases, the error rises more rapidly for the convolution-based models, while SegFormer demonstrates superior robustness. As discussed above, a different behavior is observed when considering gaussian noise, for which SegFormer already exhibits significantly higher robustness even for small corrupted regions. We believe that this advantage is largely attributed to the application of global attention mechanisms in transformers, which can effectively attend to uncorrupted areas by leveraging features from those regions or, when regions are completely perturbed, distribute robust features more efficiently across the image [13].
(a) Attack with patches (100,100)
(b) Attack with patches (200,200)
Figure 5: Accuracy in the non-corrupted region with and without (light color) localized adversarial perturbations using $\epsilon=16/255$ and $\epsilon=32/255$. The patch is always applied at the center of the image, with a size of (100, 100) in (a) and (200, 200) in (b), respectively.
(a) Segformer Attack
(b) PSPNet Attack
Figure 7: Illustration of the impact of using the Region-aware multi-attack analysis (Alg.1) on Segformer (a) and PSPNet (b) for two different images in the Cityscapes validation set. The use of multiple attacks allows for coverage of different parts of the input pixels, highlighting the complexity of solving a multi-objective adversarial optimization problem for semantic segmentation (SS) and demonstrating how the proposed algorithm can overcome this challenge. The attacks shown here use patches of size (200, 200) with $\epsilon=32/255$.
6.3 Evaluation of Adversarial Robustness
To further understand localized robustness, we evaluate the performance of the semantic segmentation (SS) models using the proposed region-aware multi-attack analysis (Algorithm 1), aiming for the identification of the worst-case scenario for spatial robustness.
As discussed in Section 5, for adversarial evaluation, it is important to note that we focus on the analysis of non-corrupted regions ($A_{\overline{M}}$) only. This is because evaluating corrupted regions under localized attacks is straightforward, as any model will typically exhibit near-zero accuracy within the corrupted area. In contrast, analyzing the non-corrupted regions provides a more meaningful assessment of spatial robustness.
In Figure 5(a) and Figure 5(b), we report the accuracy in the non-corrupted regions under attack, compared to the non-attacked case (‘clean’, shown in the lightest color in the bars), using localized adversarial attacks with patch sizes of (100, 100) and (200, 200), centered in the image, respectively. In both cases, we tested attacks with $\epsilon$ values of $16/255$ and $32/255$ (indicated by different colors in the plots). The multi-attack of Algorithm 1 is performed with $N_{\text{Att}}=3$ attacks and 50 optimization iterations (see Eq. (10)) per attack. Further analysis of these parameters is discussed in the following.
As shown in the plots, the trend is completely opposite to what we observed for natural corruptions. In this case, transformers exhibit lower robustness, primarily due to the global attention mechanisms they employ [13, 34, 33]. For example, even with a small patch size (100, 100), the accuracy drops significantly from 0.73 to 0.25 and 0.17 for $\epsilon=16/255$ and $\epsilon=32/255$, respectively.
In contrast, convolution-based models, such as DeepLabV3 with ResNet backbones, demonstrate significantly higher robustness. This robustness can be attributed to the limited local receptive field inherent to convolutional layers [21]. However, it is important to note that these models are not purely convolutional, as they also integrate attention mechanisms to enhance performance. This integration can result in substantial accuracy drops, as seen with PSPNet when using patches of size (200, 200). This observation is consistent with previous studies, which suggest that even convolutional models employing high-context exploitation through attention-based mechanisms (such as pyramidal squeezing and excitation blocks) can still be vulnerable to certain localized adversarial perturbations [29, 24].
Figure 8: Analysis of the impact of the proposed localized adversarial analysis when considering multiple attacks $N_{\textit{Att}}$ in Algorithm 1(a), with the number of iterations fixed to 50 and $\epsilon=16/255$ using patches of size (200, 200); and varying the number of optimization iterations per attack in (b), with the number of attacks fixed to 3 and the same settings for $\epsilon$ and patch size.
Benefits of Multi-Attacks
To better highlight the benefits provided by Algorithm 1, we present in Figure 8(a) the increase of the RCE, measured with respect to the clean accuracy in the non-corrupted regions, as the number of attacks $N_{\text{Att}}$ in Algorithm 1 increases. For these tests, we used patches of size (200, 200), $\epsilon=16/255$, and 50 optimization iterations per attack.
Specifically, with a single attack (#1), the multi-attack is equivalent to running a standard PGD attack over the entire targeted region. However, as the number of attacks increases, the algorithm begins focusing on fooling regions that have not yet been attacked, thereby expanding the coverage to new regions that were previously unaffected. In fact, as one may expect, the effectiveness of the multi-attack significantly improves with more attacks, as indicated by the increase in RCE. This effect is particularly pronounced in models with high attention mechanisms, such as transformers and PSPNet, demonstrating the importance of the multi-attack approach in identifying worst-case scenarios of localized adversarial perturbations. For example, the RCE of SegFormer models increases significantly, rising from approximately 0.6 with a single attack to around 0.95 when using five attacks ($N_{\text{Att}}=5$).
We also provide illustrations of the effect of multiple attacks for SegFormer and PSPNet in Figures 7(a) and 7(b). As shown in the plots, the adoption of multiple attacks with awareness of the fooling regions allows for overcoming issues in the multi-objective adversarial problem for semantic segmentation (SS). In Figure 7(a), at each attack, the untargeted multi-attacks target different classes, highlighting the inter-class problem discussed previously in Figure 2(a). Similarly, for PSPNet (Figure 7(b)), addressing different fooling regions leads to a third attack capable of affecting the entire image, thereby overcoming potential objectives that previously limited the effectiveness of earlier attacks.
Impact of the Number of Optimization Iterations
We also analyze in Figure 8(b) the impact of varying the number of optimization iterations used at line 9 in Algorithm 1, fixing the number of attacks $N_{\text{Att}}=3$. As expected, the RCE increases when using more optimization iterations, such as 200 or 250, highlighting that the complexity of the multi-objective optimization problem in this context requires a higher number of iterations to improve effectiveness. However, we acknowledge that for large-resolution datasets, as Cityscapes, a limited number of iterations (e.g., 50) is sufficient to provide meaningful and consistent comparisons, as demonstrated in the plots. This balance helps find a trade-off between computational cost and robustness evaluation, particularly when dealing with high-resolution images, which are computationally expensive due to the gradient computations required to update the attack at each step.
Figure 9: Differences in relative corruption error when applying localized adversarial attacks at the center of the image versus the bottom-left corner. Patches of size (200, 200) and $\epsilon=16/255$ are considered.
On the Position of the Attacks
Finally, we highlight the importance of the position of the adversarial patch within the image. The previous attacks were conducted with the patch placed at the center of the image, which represents the worst-case position for standard convolutional models due to its ability to affect the largest possible receptive field [21, 28, 24]. However, it is important to analyze how the effectiveness of localized adversarial attacks decreases when the patch is shifted to more constrained areas, such as the corners of the image.
In Figure 9, we compare the difference between the RCE when attacking the central area versus the one when attacking the bottom-left corner. Patches of size (200, 200) and $\epsilon=16/255$ were considered. Specifically, the larger the difference, the more the model demonstrates robustness with respect to the attack position.
As expected, transformer-based architectures, which rely on global attention and process images through patches, exhibit lower sensitivity to the position of adversarial patches. Even when the patch is moved away from the center, the model’s response remains close to the worst-case scenario. This highlights the critical risk of attacks for such architectures, being effective from multiple locations within the image, leading to the same adversarial outcome. In contrast, convolutional models exhibit more variable behavior depending on the patch’s position. However, since the tested models are not purely convolutional and incorporate attention mechanisms, their robustness is still partially affected when the patch is placed in a corner. For example, BiSeNet with a ResNet-101 backbone and PSPNet demonstrate differences of approximately 0.05 and 0.06 in RCE between central and corner positions, respectively, indicating that localized attacks can still significantly degrade their performance.
(a) $f_{1}$: DeepLab_MobileNet
$f_{2}$: SegFormer._b0
(b) $f_{1}$: DeepLab_MobileNet
$f_{2}$: SegFormer._b1
(c) $f_{1}$: DeepLab_ResNet
$f_{2}$: SegFormer._b0
(d) $f_{1}$: DeepLab_ResNet
$f_{2}$: SegFormer._b1
(e) $f_{1}$: Bisenet-R101
$f_{2}$: SegFormer._b0
(f) $f_{1}$: Bisenet-R101
$f_{2}$: SegFormer._b1
(g) $f_{1}$: DDRNet23
$f_{2}$: SegFormer._b0
(h) $f_{1}$: DDRNet23
$f_{2}$: SegFormer._b1
Figure 11: Analysis of the natural and adversarial errors, computed as discussed in Equations 12 and 13, respectively, and clean accuracy when considering different models as $f_{1}$ and $f_{2}$ in the ensemble strategy defined in Equation 11, for different configurations of $\gamma$. The results aim to explore the existence of a possible trade-off to balance robustness against localized natural corruptions and adversarial perturbations.
6.4 Searching for a Trade-off in Ensembling
Given the results obtained in the above analyses of natural perturbations and adversarial attacks, we acknowledge that transformer architectures generally exhibit greater robustness to localized natural corruptions by effectively reducing misclassifications, not only within the corrupted regions but also in the surrounding non-corrupted areas. However, their reliance on high attention mechanisms makes them more vulnerable to localized adversarial attacks, resulting in a significant drop in accuracy compared to convolutional models.
Following this observation, we emphasize the importance of searching for a trade-off that balances natural robustness and worst-case adversarial robustness, which could be particularly beneficial in safety-critical domains such as autonomous driving. To address this challenge, we investigate the use of model ensembling, with natural and adversarial localized robustness as key metrics to evaluate the effectiveness of the ensemble strategy. To explore this approach, we first define a test-time weighted ensemble strategy [12] as:
$g_{\gamma}(f_{1},f_{2})=\gamma\cdot f_{1}+(1-\gamma)\cdot f_{2},$ (11)
where $f_{1}$ and $f_{2}$ are two models selected for the ensemble, and $\gamma$ is a parameter that balances the importance of the softmax scores of model $f_{1}$ with respect to $f_{2}$ in the final model $g$.
Specifically, for our tests, we analyze pairs of models where $f_{1}$ is a model with higher robustness to localized adversarial attacks (e.g., DeepLab-ResNet101, DeepLab-MobileNet, BiSeNet-ResNet101, and DDRNet23), and $f_{2}$ is demonstrates strong robustness to natural corruptions but lower robustness to localized adversarial attacks (e.g., SegFormer_b0 or SegFormer_b1).
To understand the benefits of ensembling, we redefine the RCE for natural and adversarial perturbations, as follows, to account for the combination of $f_{1}$ and $f_{2}$. Most important to us are the RCE for natural perturbations within the corrupted area, i.e.,
$\textit{CE}_{\textit{Nat}}(g,f_{2},C,M)=\frac{A_{\mathcal{P}}(f_{2},C,M)-A_{\mathcal{P}}(g,C,M)}{A_{\mathcal{P}}(f_{2},C,M)},$ (12)
where $\mathcal{P}_{i}(M)=M_{i}/|M|$ and $C$ is a set of naural corruptions, and the RCE for adversarial perturbations outside the corrupted area, i.e.,
$\quad\textit{CE}_{\textit{Adv}}(g,f_{1},C,M)=\frac{A_{\mathcal{P}}(f_{1},C,M)-A_{\mathcal{P}}(g,C,M)}{A_{\mathcal{P}}(f_{1},C,M)},$ (13)
where $\mathcal{P}_{i}(M)=\overline{M}_{i}/|\overline{M}|$ and $C$ is a set of adversarial corruptions.
The rationale behind these error definitions is that, since $f_{2}$ and $f_{1}$ are chosen as the best-performing models for natural and adversarial robustness, respectively, this formulation allows us to understand how varying $\gamma$ influences improvements in natural robustness and the corresponding drop in adversarial robustness.
The results of Figure 11 confirm the expected trend. As we increase $\gamma$, the ensemble model becomes more robust to natural perturbations but also increasingly susceptible to localized adversarial attacks. Note that, in this case, the attacks are performed directly on the ensemble model, considering the corresponding $\gamma$ value. Interestingly, for certain values (e.g., $\gamma=0.4$), the adversarial robustness is largely preserved, while we observe a significant reduction in the natural error. This phenomenon is particularly noticeable when DeepLab architectures are used as $f_{1}$, as the natural error decreases slightly before the adversarial error curve begins to rise. This demonstrates that a positive trade-off between natural and adversarial robustness can be found.
Additionally, we report the mean accuracy of the ensemble on clean inputs in the top portions of the plots to ensure fairness and confirm that there is no drop in clean accuracy. The error of clean accuracy is computed with respect to the best-performing model between $f_{1}$ and $f_{2}$. As shown in the plots, the use of a well-selected $\gamma$ gets also to negative errors, indicating that the ensemble strategy improves performance on the clean validation set.
We believe this analysis highlights the importance of rethinking model ensembling to achieve results beyond raw performance. This approach can be particularly effective when combining architectures with orthogonal strengths, such as robustness to localized natural corruptions and robustness to localized adversarial attacks.
7 Discussion and Conclusion
This work presented an extensive analysis of the robustness of segmentation models under localized corruptions, considering both natural and adversarial attacks. We first introduced metrics that extend the classic pixel-wise accuracy—commonly used in segmentation—to provide a more comprehensive analysis of spatial robustness, both within and beyond the corrupted regions. We then proposed a framework that enables a systematic and efficient evaluation of localized natural and adversarial corruptions by allowing the selection of key parameters, including corruption ratios and model patch sizes.
Our analysis of perturbations also improves upon existing strategies in the literature by demonstrating that, in dense prediction tasks such as semantic segmentation, targeting multiple pixels introduces a complex multi-objective problem. This complexity makes it difficult to assess the worst-case adversarial scenario using a single localized perturbation. To address this, we proposed region-aware multi-attack adversarial analysis, a method that applies multiple attacks while continuously adjusting the fooling area, thus providing a more realistic assessment of model robustness across different parts of the output.
All findings were evaluated across a diverse set of segmentation models in a driving scenario, specifically using the Cityscapes dataset [8]. Our results revealed distinct differences in model behavior under natural and adversarial corruptions. Furthermore, we highlighted the importance of finding a trade-off between these two robustness aspects, which are not necessarily correlated. As a preliminary step toward addressing this challenge, we explored ensemble strategies that combine convolutional-based models with transformer-based architectures.
For future work, we aim to deepen the analysis of such a trade-off also considering other tasks and theoretical insights. Another important direction is integrating the proposed localized robustness analysis into model training strategies, particularly through the lens of localized augmentation techniques (e.g., Cutout[11]) and CutMix[38], which are not commonly applied in complex scene understanding tasks like semantic segmentation. In this work, we focused on standalone models trained with conventional image-level augmentations, leaving this integration as a promising avenue for further research.
To conclude, our study represents an important advancement in understanding the spatial robustness of segmentation models, while also identifying key challenges and open research directions for future investigations.