### v3

# Benchmarking the Spatial Robustness of DNNs via Natural and Adversarial Localized Corruptions

Abstract
The robustness of deep neural networks is a crucial factor in safety-critical applications, particularly in complex and dynamic environments (e.g., medical or driving scenarios) where localized corruptions can arise. While previous studies have evaluated the robustness of semantic segmentation (SS) models under whole-image natural or adversarial corruptions, a comprehensive investigation into the spatial robustness of dense vision models under localized corruptions remains underexplored. This paper fills this gap by introducing novel, region-aware metrics for benchmarking the spatial robustness of segmentation models, along with an evaluation framework to assess the impact of natural localized corruptions. Furthermore, it uncovers the inherent complexity of evaluating worst-case spatial robustness using only a single localized adversarial attack. To address this, the work proposes a region-aware multi-attack adversarial analysis to systematically assess model robustness across specific image regions. The proposed metrics and analysis were exploited to evaluate 14 segmentation models in driving scenarios, uncovering key insights into the effects of localized corruption in both natural and adversarial forms. The results reveal that models respond to these two types of threats differently; for instance, transformer-based segmentation models demonstrate notable robustness to localized natural corruptions but are highly vulnerable to adversarial ones, and vice versa for CNN-based models. Consequently, we also address the challenge of balancing robustness to both natural and adversarial localized corruptions by means of ensemble models, thereby achieving a broader threat coverage and improved reliability for dense vision tasks.
1 Introduction
In recent years, deep neural networks (DNNs) have demonstrated remarkable performance across various vision applications, increasing their potential applicability in safety-critical domains such as autonomous systems[survey_liu2024survey, PR_QIAN2022108796]. Among these scenarios, dense prediction tasks like semantic segmentation (SS) have garnered significant attention. These tasks require models to understand the semantic meaning of each pixel in a given scene, often necessitating the extraction of deep contextual information. In this context, it is crucial to assess the robustness and trustworthiness of SS models, as complex and dynamic environments can give rise to unexpected input variations that challenge the entire system reliability.
The complexity of DNNs and real-world application scenarios, as in autonomous and assisted driving, has made ensuring and evaluating robustness one of the most challenging problems in modern AI [Carlini017, biggio2018wild, Szegedy14, survey_trust]. In the literature, robustness evaluation typically involves analyzing model performance under corrupted input conditions, with two main types of corruption commonly studied: (i) Natural corruptions, which simulate faults or real-world environmental conditions such as sensor noise or adverse weather (e.g., snow, rain) [hendrycks2018benchmarking]; (ii) Adversarial corruptions (adversarial attacks), which involve intentionally crafted perturbations designed to manipulate model predictions. While adversarial attacks are often studied as a security threat, they are also a valuable tool for evaluating a model’s robustness under a close approximation of worst-case scenarios, contributing to a sense of verifiability [brau2022minimal, survey_trust].
In recent years, numerous studies have evaluated the robustness of vision models against both natural and adversarial perturbations [hendrycks2018benchmarking, Szegedy14, kamann2020benchmarking]. However, a comprehensive characterization of their behavior under localized corruptions [patch_object, rossolini_tnnls_2023, saha_role_2020], which affect only parts of the image, remains insufficiently explored, particularly for dense prediction tasks such as semantic segmentation. Indeed, although poorly explored, localized perturbations are highly relevant to real-world scenarios (e.g., sunlight affecting specific regions of an input image) and need to gather more attention in robustness studies. Exploring localized corruptions allows addressing critical questions about the spatial robustness of semantic segmentation models: To what extent do corrupted regions influence predictions in both the corrupted and uncorrupted areas of an image? How effectively can models leverage uncorrupted regions to mitigate the impact of corrupted ones? Addressing these questions is crucial for advancing the robustness and reliability of scene-understanding models in safety-critical environments.
This paper addresses the above challenges by studying the effect of localized corruptions across different models through the lens of both natural transformations and adversarial perturbations. In particular, the work begins by formalizing the concept of localized corruptions and proposing ad-hoc metrics specifically designed to evaluate spatial robustness in different regions of the input for SS models.
An analysis framework for benchmarking the spatial robustness of SS models against localized natural corruptions is then introduced, enhancing the understanding of the reliability of DNNs in practical scenarios, such as in driving environments, where localized corruptions can occur.
Next, critical aspects of adversarial attacks within the context of SS tasks are investigated, highlighting that identifying a single localized perturbation capable of comprehensively evaluating spatial robustness (e.g., misclassifying as many pixel predictions as possible) is particularly challenging. This is due to the multi-objective nature of the adversarial optimization problem. To address this challenge, we propose the region-aware multi-attack adversarial analysis, a novel evaluation algorithm that iteratively generates localized perturbations, each targeting regions previously unaffected by prior attacks. The idea behind the proposed algorithm builds on the observation that regions initially considered robust (because they are not misclassified by a single localized attack) can, instead, be compromised when repeated runs of the same attack are applied with more targeted configurations. This enables a clearer understanding and evaluation of worst-case localized perturbations, revealing vulnerabilities that may not be sufficiently captured by standard localized attacks.
The proposed framework, considering both localized natural and adversarial corruptions, is then applied in an experimental evaluation to benchmark the spatial robustness of $14$ SS models in driving scenarios [cordts2016cityscapes]. Most interestingly, we observed that addressed convolution-based models are significantly vulnerable to localized natural corruptions, while addressed transformer-based models exhibit higher robustness. Conversely, the opposite trend arises when considering localized adversarial corruptions: convolution-based models appear more robust, whereas transformer-based models appear highly vulnerable due to the heavy exploitation of global attention mechanisms. These findings raise contrastive concerns about the reliability of SS models when evaluated under worst-case safety-related scenarios. To address this, we finally investigate a promising ensembling analysis aimed at achieving a better trade-off in spatial robustness between natural and adversarial perturbations.
Overall, this study marks a significant step toward understanding localized corruptions and paves the way for developing more robust SS models in practice. [^1] To support reproducibility, we provide the code and benchmark implementation used in our experiments at the following repository: https://github.com/GiuliaMP/SpatialRobustnessBench.
In summary, this paper makes the following contributions:
  • It formalizes new metrics to evaluate the accuracy of SS models while accounting for their spatial robustness under localized corruptions. Furthermore, it extends existing approaches to study localized corruption by proposing a flexible framework for evaluating the model spatial robustness.
  • It uncovers important insights and challenges in the study of localized adversarial attacks in SS models and introduces a new evaluation method that provides a more thorough assessment of robustness by iteratively testing multiple attacks.
  • The proposed metrics, frameworks, and methods are exploited to analyze the spatial robustness of 14 SS models on the Cityscape dataset, under natural and adversarial localized perturbations, revealing contrasting behaviors between the two types of corruptions. Finally, the trade-off between natural and adversarial robustness is explored by means of ensemble analysis.
The rest of the paper is structured as follows: Section 2 discusses related work and remarks the importance of studying the spatial robustness of dense prediction models; Section 3 introduces the proposed analysis and formalizes the metrics we used; Section 4 defines the evaluation framework adopted for assessing models under localized natural corruptions; Section 5 focuses on localized adversarial perturbations; Section 6 discusses our experimental evaluations and ensemble models, and finally, Section 7 states the conclusions and discussion.
2 Related Work
Analysis and benchmarks on DNNs robustness
Deep learning models are widely used across various applications, including safety-critical domains such as autonomous driving. Given their increasing deployment in such scenarios, assessing their robustness is essential to evaluate their reliability and suitability for real-world use. To address these challenges, several benchmarks and robustness analyses have been proposed over the years to assess model reliability against both natural and adversarial corruptions. Notable examples include [hendrycks2018benchmarking, croce2022interplay, kamann2020benchmarking, schiappa2024robustness].
Regarding natural corruptions, the objective is to evaluate models’ robustness against perturbations that simulate adverse or potentially out-of-distribution conditions that can be encountered in real-world environments, such as noise, rain, or snow. One of the earliest and most influential studies in this area was conducted by Hendrycks and Dietterich, who introduced a benchmark to assess the resilience of image classification models. This benchmark comprises various natural corruptions, including weather-related disturbances and image formatting artifacts [hendrycks2018benchmarking]. Since then, these corruptions have been widely adopted and extended in subsequent studies to systematically evaluate and compare the robustness of different models [gu2022vision, bhojanapalli2021understanding, naseer2021intriguing]. For instance, Bhojanapalli et al. [bhojanapalli2021understanding] investigated the robustness of ViTs comparing it to the robustness of a ResNet-50 model under image corruptions, concluding that ViTs models are at least as robust as the CNN taking into consideration on the majority of the corruptions. While the majority of studies have focused on classification tasks, fewer works have explored the robustness of models in dense prediction tasks, such as semantic segmentation [schiappa2024robustness, kamann2020benchmarking], which have highlighted the need for more comprehensive evaluations in this new domain.
Beyond natural corruptions, other works have examined adversarial corruptions [Szegedy14, biggio2018wild], which represent a worst-case scenario from a security perspective, where perturbations are deliberately crafted to deceive models. Similar to natural corruptions, several studies have compared the robustness of different architectures against adversarial perturbations [bhojanapalli2021understanding, naseer2021intriguing, mahmood2021robustness]. Specifically, research on adversarial robustness in semantic segmentation [yuan2024towards, arnab_robustness_nodate, kamann2020benchmarking] has underscored the importance of establishing more rigorous benchmarks and developing more effective defense strategies to enhance models’ robustness in real-world applications.
Evaluation of localized corruptions and spatial robustness
While most of the aforementioned studies primarily focus on corruptions affecting the entire input, it is crucial to consider the problem of spatial robustness [saha_role_2020, rossolini_tnnls_2023, gu2022vision]. This aspect is particularly relevant for dense prediction tasks, where model predictions may be influenced by regions beyond the immediate locality of the target area. For instance, since semantic segmentation performance is usually evaluated at a pixel level, defining a robustness analysis for such a task opens the possibility to study the impact of perturbed regions on clean regions, and vice versa.
Despite its importance, research on spatial robustness under natural corruptions remains limited. For example, [gu2022vision] analyzes the robustness of ViTs compared to CNNs under natural and adversarial corruptions, concluding that ViTs generally exhibit greater resilience to natural corruptions than CNNs. However, does not specifically address dense prediction tasks. Conversely, most existing studies on spatial robustness have focused on physically-realizable adversarial settings, e.g., adversarial patches [pintor_imagenetpatch, brown_adversarial_2018, liu2025_patch, metzen2021meta, rossolini_iccps24], while only a few works have explored the effects of single localized adversarial perturbations [nakka_indirect_2020], which instead reveal intrinsic and human-imperceptible weaknesses of the model. Additionally, there is still a lack of comparative analysis on transformer-based models in both cases.
Differently from these works, this paper shows that, from an adversarial perspective, a single attack is not sufficient to properly represent the true spatial robustness of dense prediction models. To this end, it proposes an ad hoc evaluation analysis that iteratively adjusts the fooling regions across multiple attacks in the same area, enabling the assessment of regions that conventional attacks might leave unaffected. Furthermore, to address the lack of analysis on natural localized robustness, an aspect particularly relevant for safety assessments, this work introduces a schematic framework and ad hoc metrics to benchmark models in these settings.
To the best of our knowledge, this work provides the first comprehensive analysis of spatial robustness in the context of both natural and adversarial perturbations for semantic segmentation. This study allowed uncovering important experimental insights when comparing transformer-based and convolution-based architectures and proposing a promising ensembling strategy.
3 Spatial Robustness Analysis
This section first provides some preliminary and background information to theoretically define localized corruptions within the context of SS models. Then, new metrics for spatial robustness are proposed to generalize the classic pixel-wise accuracy score used in semantic segmentation.
3.1 Background and Formalisms
Preliminaries
We consider input images of size $H\times W$ and denote by $\mathcal{I}$ the set of the image’s pixels, which are then individually referred to using the index $i$. A semantic segmentation model designed to classify each pixel into one of $N$ classes is represented by a function $f:x\mapsto\mathbb{R}^{(H\cdot W)\times N}$, which outputs a per-class distribution of scores for each pixel $i$. The vector of output scores for each pixel $i$ is denoted by $f_{i}$ while its component related to the $j$-th class is denoted by $f^{j}_{i}$. The semantic label predicted for pixel $i$ is determined by selecting the class with the highest score for that pixel, i.e., $\operatorname*{argmax}_{j\in\{1,\dots,N\}}f_{i}^{j}(x).$ Finally, the corresponding ground-truth segmentation map for the image $x$ is denoted by $y\in\mathbb{N}^{H\times W}$, where each pixel is assigned its true class label, i.e., $y_{i}\in\{1,\dots,N\}$.
Corruption Areas and Ratio
To test the spatial robustness of SS models, we consider corruptions applied to only part of the image. Specifically, following the notation adopted in previous work [nakka_indirect_2020, rossolini_tnnls_2023], we define a corrupted area $M$ as a binary mask in the input space $M\in\{0,1\}^{H\times W}$, which indicates where the corruption is applied. For any 2D pixel with index $i$, $M_{i}=1$ denotes the pixel is corrupted, while $M_{i}=0$ denotes that it is not. The corrupted image $x_{M}^{c}$, given a mask $M$ and a corruption function $c(\cdot)$, is computed as:
$x^{c}_{M}=c(x)\cdot M+x\cdot(1-M),$ (1)
where $c(\cdot)$ can represent either a natural transformation or an adversarial attack. Additionally, we introduce the ratio of the perturbed area as $r=S/(H\cdot W)$, where $\mathcal{S}$ is the number of corrupted pixels, computed from $M$ as $\mathcal{S}=\sum_{i=1}^{H\cdot W}M_{i}$. It follows that the number $\overline{\mathcal{S}}$ of non-corrupted pixels is $\overline{\mathcal{S}}=\sum_{i=1}^{H\cdot W}1-M_{i}$.
3.2 Metrics for Spatial Robustness
We define metrics for studying the effect of localized corruptions on predictions in different areas of the image, by generalizing both the classic accuracy metrics used for SS and those applied for testing the robustness of image-classification DNNs against corruptions.
Classic Corruption Accuracy
Deriving from the analysis of [hendrycks2018benchmarking] in image classification, given a model $f$ and a set of possible corruptions $\mathcal{C}$ (e.g., Gaussian noise, adversarial perturbation or $\{\emptyset\}$ to denote no corruptions), the robustness accuracy of an SS model on a dataset $\mathcal{D}$ is defined as:
$A(f,\mathcal{C})=\frac{1}{|\mathcal{C}|\cdot|\mathcal{I}|\cdot|\mathcal{D}|}\sum_{c\in\mathcal{C}}\sum_{(x,y)\in\mathcal{D}}\sum_{i\in\mathcal{I}}\mathds{1}\big(f_{i}(c(x))=y_{i}\big),$ (2)
where $|\mathcal{C}|$ is the number of corruption types, $|\mathcal{I}|$ is the number of pixels in the input, $|\mathcal{D}|$ is the size of the dataset, and $\mathds{1}(\cdot)$ is an operator that returns ‘1’ if a condition is ‘True’, otherwise ‘0’. Following this metric, we can also define the Corruption Error as:
$\textit{CE}(f,\mathcal{C})=\frac{A_{\textit{base}}-A(f,\mathcal{C})}{A_{\textit{base}}},$ (3)
where $A_{\textit{base}}$ refers to the accuracy of a reference model (either $f$ or another one, e.g., AlexNet, as used in [hendrycks2018benchmarking]). This enables a better understanding of the impact of a given corruption with respect to a reference value.
Localized Corruption Accuracy
We extend the analysis of the accuracy to consider the case of localized corruption, identified by a corruption mask $M$. The accuracy of a model $f$ under a corruption $c$ applied in the areas denoted by $M$ is defined as:
$A_{\mathcal{P}}(f,\mathcal{C},M)=\frac{1}{|\mathcal{C}|\cdot|\mathcal{D}|}\sum_{c\in\mathcal{C}}\sum_{(x,y)\in\mathcal{D}}\mathcal{K}_{\mathcal{P}}(f,x,c,M,y),$ (4)
with
$\mathcal{K}_{\mathcal{P}}(f,x,c,M,y)=\sum_{i\in\mathcal{I}}\mathds{1}\big(f_{i}(c(x))=y_{i}\big)\cdot\mathcal{P}_{i}(M),$ (5)
where $\mathcal{P}_{i}(M)$ is a spatial importance function designed to weigh the significance of misclassifying a pixel $i$ based on the corruption mask $M$. This function plays a critical role in assessing the impact of regions of the input image, allowing the analysis to either focus on regions of higher relevance or those affected by corruption. While, in general, spatial importance functions can be arbitrarily defined, in this study we focus on two specific definitions:
  • Non-Corrupted Region Analysis ($A_{\overline{M}}$): this approach focuses on evaluating the spatial robustness of the model by analyzing regions unaffected by corruptions, uniformly weighting all non-corrupted pixels, i.e. $\mathcal{P}_{i}(M)=(1-M_{i})/\overline{\mathcal{S}}$.
  • Corrupted Region Analysis ($A_{M}$): this approach is the dual of the one above, focusing on corrupted pixels only, i.e., $\mathcal{P}_{i}(M)=M_{i}/\mathcal{S}$.
Notice that both $A_{\overline{M}}$ and $A_{M}$ are normalized by the number of pixels in their respective regions, ensuring fairness in evaluating and comparing each region. Such localized metrics reveal whether the corruption affects predictions beyond its regions, or whether the model can leverage clean informations from outside the perturbed areas to correctly classify the perturbed pixels.
The first definition ($A_{\overline{M}}$) is particularly useful for assessing how well the model remains accurate in non-corrupted regions, offering insights into the capability of corruptions to extend beyond the actual corrupted region. Conversely, the second definition ($A_{M}$) focuses on the model’s ability to recover accurate predictions in corrupted regions, potentially leveraging information from non-corrupted regions. This analysis is more meaningful when studying natural corruptions, as adversarial attacks can easily lead to complete misclassification of corrupted regions [rossolini_tnnls_2023, nakka_indirect_2020], making robustness assessments in such areas less informative. It is worth noting that the metrics presented above can be computed for multiple corruption regions: different settings are considered in our experimental evaluation (Section 6). For localized adversarial perturbations, we follow traditional approaches that focus on statically targeting specific areas of the image, such as the center or a corner (e.g.,  [rossolini_tnnls_2023, nakka_indirect_2020]). The rationale behind this choice is that these positions allow evaluating the worst- or best-case scenarios of a model’s receptive field [luo2016understanding].
Different considerations can be made when assessing the impact of arbitrarily-placed localized corruptions. For natural corruptions, we consider the above metrics while selecting the corruption regions randomly, leading to the following formulation:
$A_{\mathcal{P}}(f,\mathcal{C})=\frac{1}{|\mathcal{C}|\cdot|\mathcal{D}|}\sum_{c\in\mathcal{C}}\sum_{(x,y)\in\mathcal{D}}\mathcal{K}_{\mathcal{P}}(f,x,c,M_{r}(x),y),$ (6)
where $M_{r}(x)$ is a mask that determines the corrupted regions, with patches randomly selected for each input $x$ while respecting the ratio $r=S/(H\cdot W)$ of the corrupted area.
It is finally important to remark that all the proposed metrics focus on pixel-wise accuracy for semantic segmentation. Other metrics, such as the mean intersection over union (MIoU), are also frequently used in the field of SS and our analysis can be extended to support them. For instance, no significant modifications are required to support MIoU.
4 Localized Natural Corruptions
Figure 1: Illustrations of the proposed framework for evaluating localized natural corruptions. On the left, two example configurations define the settings for localized corruption analysis: patch size, corruption ratio, corruption type, and the severity level. The framework returns a dataset sample as a tuple consisting of the original image, the ground truth label, the corrupted image, and a mask highlighting the corrupted areas.
This section presents the framework designed to evaluate localized natural corruptions. The framework includes all the corruption types available in ImageNet-C [hendrycks2018benchmarking] which, however, does not support localized corruptions. In particular, we consider different transformations $c$, which can be applied at various severity levels $s$ (i.e., 1, 2, 3, 4, 5), as selected by the user. A natural corruption applied to an input $x$ with severity level $s$ is denoted by $c(x)=\Gamma(x,s)$, where $\Gamma$ represents a natural transformation, such as gaussian noise, gaussian blur, motion blur, etc.
To enable a systematic, reproducible, and broad evaluation of spatial robustness, the corrupted region indicated by $M$ is automatically selected according to the following pipeline. The image $x$ is first divided into non-overlapping patches with size $\Delta=(\Delta_{x},\Delta_{y})$, thereby obtaining a total of $P=P_{x}\times P_{y}$ patches with $P_{x}=\lceil W/\Delta_{x}\rceil$, $P_{y}=\lceil H/\Delta_{y}\rceil$, where $H$ and $W$ are the height and width of $x$, respectively. Given a specified corruption ratio $r$, each patch, and therefore all its pixels, is selected to be perturbed with probability $r$. All pixels within a selected patch are assigned a value of $1$ in $M$, while the others have a value of 0.
In Figure 1, we present an illustration of a possible configuration used within the tested framework and the output of a dataset sample in the corrupted validation set of Cityscapes [cordts2016cityscapes]. Specifically, a sample corresponds to a tuple containing the original image and its label (kept available to allow comparisons during the evaluation), along with the perturbed image and the corruption region mask $M$. During evaluation, for each image, the set of patches to be perturbed (and thus the mask $M$) is recomputed, effectively randomizing the selection of the perturbed areas for the evaluation.
5 Localized Adversarial Corruptions
This section presents our novel analysis for localized adversarial perturbations. Although previous work already addressed the study of localized adversarial perturbations against SS models [nakka_indirect_2020], we next highlight the difficulties of performing a comprehensive spatial adversarial evaluation using a single attack. To this end, we begin by discussing classic approaches and pointing out their drawbacks. Finally, we introduce our region-aware multi-attack algorithm designed to address these issues and offer a better understanding of the worst-case spatial robustness of computer vision models.
5.1 Attacks for SS from previous work
In the context of classic adversarial perturbations [Szegedy14, biggio2018wild], the main idea is to craft a specific adversarial noise $\delta$ that realizes a worst-case perturbation constrained by a magnitude $\varepsilon$ under a specific norm, e.g., the $l_{\infty}$ norm, such that $\|\delta\|_{\infty}\leq\varepsilon$. Adversarial perturbations are typically computed by optimizing an adversarial loss function $\mathcal{L}$, e.g., cross-entropy in image classification, which aims to find a perturbation $\delta$ for an input $x$ that forces the model to make incorrect predictions:
$\delta=\arg\max_{\|\delta\|_{\infty}\leq\varepsilon}\mathcal{L}(f(x+\delta),y),$ (7)
Recalling Equation 1, it is possible to express inputs affected by localized attacks with a perturbation $\delta$ and a mask $M$ as follows:
$x_{M}^{\delta}=(x+\delta)\cdot M+x\cdot(1-M).$ (8)
When interested in dense prediction tasks, such as semantic segmentation, where predictions are made at the pixel level, the perturbation is generally computed to maximize the misclassification of all or a subset of pixels in the image [nakka_indirect_2020, rossolini_tnnls_2023]. In the latter case, it is possible to define a fooling region as a subset of the input space for which we intend to induce a misprediction. The fooling region can be arbitrarily defined by a mask $F\in\{0,1\}^{H\times W}$ that does not necessarily need to overlap with the corruption mask $M$. In this context, the perturbation $\delta$ can be optimized as follows:
$\delta=\arg\max_{\|\delta\|_{\infty}\leq\varepsilon}\mathcal{L}_{\mathcal{F}}(f(x^{\delta}_{M}),y),$ (9)
where $\mathcal{L}_{\mathcal{F}}$ is an adversarial loss designed to focus the model performance in the fooling region $F$ only. In practice, to solve the above optimization problem, the perturbation $\delta$ can be iteratively updated with step size $\alpha$ to maximize the classification error [nakka_indirect_2020, pgd_attack]:
$\delta_{i+1}\leftarrow\delta_{i}+\alpha\cdot\text{sign}\left(\nabla_{\delta}\mathcal{L}_{\mathcal{F}}\right).$ (10)
5.2 Limitations of previous work
Although previous approaches are capable of crafting strong perturbations by corrupting as many pixels as possible, we argue that comprehensively evaluating the spatial robustness of semantic segmentation models requires multiple attacks applied to the same area. A single perturbation, even when computed using a strong attack, cannot cover all regions susceptible to misprediction. This happens because, as shown in Equation 9, crafting an optimal localized adversarial perturbation involves solving a complex multi-objective optimization problem, where the goal is to increase the loss for each pixel in the image, representing a wide range of distinct objectives. Many of these objectives are difficult to solve simultaneously, often leading to suboptimal or locally optimal solutions. To support this observation, we provide representative analysis results in Figures 2(a) and  2(b) (discussed in the next paragraphs), which highlight, from both inter-class and intra-class perspectives, that adversarial attacks targeting specific subsets of pixels may not generalize effectively to others, making it particularly challenging to craft a single localized perturbation that generates worst-case adversarial effects.
(a) Inter-class target attack analysis.
(b) Intra-class target attack analysis
Figure 2: Analysis of limitations for single localized perturbations in $S$: (a) Inter-class analysis: clean prediction, attacked image, and predictions across different target classes, shown in sequence. (b) Intra-class analysis: Targeted multi-attack on the class ‘Car’. The first row displays the fooling area, the attacked image, and the cumulative adversarial output. The second row shows the output after each attack iteration, while the last row visualizes how the fooling region is selected and which pixels are actually affected at each attack.
Let us first focus on inter-class attacks with reference to Figure 2(a), where localized perturbations are applied to the center of the image ($100\times 100$, using iterative targeted attacks [nakka_indirect_2020] with $\epsilon=32/255$). In this example, attacks are performed each targeting different classes in separate runs (‘road’, ‘sky’, and ‘car’). The outputs of the addressed targeted attacks and their corresponding classes are shown from the third to the fifth subfigures. Additionally, we also show the output prediction of an untargeted attack (targeting all classes), in the last subfigure. The results demonstrate that the adversarial patterns produce distinct adversarial effects. This highlights that the multi-objective optimization landscape for different target classes is highly divergent, evidencing difficulties for a targeted attack to affect the areas interested by others. Furthermore, when observing the output of the untargeted attack, it closely resembles the output of the ‘road’ attack, suggesting a bias in the multi-objective optimization process, possibly given by the fact that the majority of pixels of the image are classified as ‘road’.
We then analyze the same problem in the context of intra-class attacks. In this case, as shown in Figure 2(b), we crafted attacks to misclassify pixels belonging to the ‘car’ class. We started with an initial fooling region, highlighted in the first subfigure, consisting of all pixels labeled as ‘car’ in the ground truth. After executing the attack of Equation 9 a first time (denoted by ‘attack 1’) using the initial fooling region, we were capable of misclassifying a limited subset of the targeted pixels only (note the attack output in the second row and the corresponding red pixels in the third row of Figure 2(b)). To carry out a second attack (attack 2), we excluded from the initial fooling area the pixels that were successfully attacked in the previous attack. This adjustment allows us to investigate whether previously attacked pixels negatively influence the optimization process for the remaining pixels classified as ‘car’. The restricted fooling region for the second attack is shown in the third row. As can be observed from Figure 2(b), the second attack expands the attacked area, indicating that previously attacked regions can indeed hinder the optimization process when exploring new attackable regions. This process is repeated for a third attack (last columns in the second and third rows of the figure), confirming the previous observation.
Finally, for completeness, we present the cumulative adversarial output in the third subfigure (first row, third column). This cumulative output assigns each pixel its adversarial misclassified prediction if at least one attempted attack has misclassified it (the first misclassification is considered for simplicity); otherwise, it retains the original model prediction. This approach, provides a clearer worst-case understanding of whether each given pixel can be misclassified by at least one attack and will be used in the algorithm proposed in the next subsection.
5.3 Achieving Stronger Spatial Adversarial Analysis
The observations made in the previous section allow concluding that classically applied single instances of SS attacks provide an incomplete assessment of the worst-case spatial adversarial robustness of a model. To address this limitation and assess a more comprehensive benchmark under localized adversarial perturbations, a method called region-aware multi-attack adversarial analysis is proposed in Algorithm 1. The algorithm builds upon the observations made above and is next explained step-by-step. At a high level, the method iteratively generates a new localized adversarial attack on a region $M$ while progressively refining the fooling region $\mathcal{F}$ at each attack run, thereby maximizing the coverage of attacked pixels.
Algorithm 1 Region-aware Multi-Attack Adversarial Analysis
1:
2:Model $f$; input $x$; labels $y$; Num. of attacks $N_{\text{Att}}$
[…]
7:for $i\leftarrow 1$ to $N_{\text{Att}}$ do
8:  $\mathcal{F}\leftarrow\mathcal{F}\cdot\Psi_{\text{corr}}$ $\triangleright$ Update fooling region
9:  $x_{\text{adv}}\leftarrow\textit{Attack}_{M}(x,y,f,\mathcal{F})$
10:  $\hat{y}_{\text{adv}}\leftarrow\text{argmax}(f(x_{\text{adv}}))$ $\triangleright$ Preds after attack
11:  $\Psi_{\text{cur}}\leftarrow(\hat{y}_{\text{adv}}=y)$ $\triangleright$ Current correct pixels
12:  $\bar{\Psi}\leftarrow(1-\Psi_{\text{cur}})\cdot\Psi_{\text{corr}}$ $\triangleright$ Newly misclassified pixels
13:  $\Psi_{\text{corr}}\leftarrow\Psi_{\text{cur}}\cdot\Psi_{\text{corr}}$ $\triangleright$ Update correct aux. mask
14:  $\hat{y}_{\text{O}}\leftarrow\Psi_{\text{corr}}\cdot\hat{y}_{\text{O}}+\bar{\Psi}\cdot\hat{y}_{\text{adv}}$ $\triangleright$ Update cumulative output
15:end for
16:return $\textit{Eval}(\hat{y}_{\text{O}},y)$
The proposed algorithm takes as input a given image $x$, its labels $y$, and an SS model $f$, and aims to evaluate the robustness of the image by running multiple attacks ($N_{\textit{Att}}$), applied to the same area $M$, based on a given method (referred to as $\textit{Attack}_{M}$ in the algorithm). At the beginning of the algorithm, the cumulative output is initialized to the correct model predictions, and the initial fooling region consists of all pixels that are correctly predicted in the non-perturbed output (line 4) [^2] We address the untargeted case for simplicity and to generalize as much as possible, covering as many pixels as possible. However, the approach can be easily extended to a targeted formulation.. Subsequently, the attack iteratively targets the pixels in the fooling region using a classic iterative attack (e.g., Equation 10, line 7), computes the corresponding output (line 8), and updates the fooling region by identifying pixels in the preceding fooling region that have not yet been misclassified in the current attack (lines 9–12 and line 6). At the end of each attack, the cumulative adversarial output is updated to replace the predictions of newly misclassified pixels that had previously been correctly classified in earlier attacks.
Finally, the cumulative output is evaluated according to the desired metrics (Eval in the algorithm). It is important to remark that the cumulative output is not derived from the model’s prediction on a single perturbed sample but rather from the cumulative worst-case attacked outputs, obtained by aggregating misclassifications across multiple localized perturbations, all concentrated in the same area. From a safety perspective, this approach provides a stronger and more accurate understanding of the worst-case robustness of each pixel, demonstrating the existence of at least one perturbation capable of misclassifying them.
Please note that, while the proposed formulation is entirely untargeted, this strategy also helps mitigate intra-class issues that may arise even in targeted attacks, as discussed above and illustrated in Figure 2(b).
6 Evaluation of Spatial Robustness
In the following, we first describe the experimental setup. Then, we evaluate the spatial robustness, based on the proposed metrics, of several models against localized natural corruptions and localized adversarial attacks separately. Finally, we discuss the importance of addressing these two aspects together and propose an ensemble-based analysis aimed at balancing localized natural and adversarial robustness.
6.1 Experimental Setup
The experiments were conducted with NVIDIA A100 GPUs, using the validation set of Cityscapes [DBLP:conf/cvpr/CordtsORREBFRS16], which serves as the reference driving dataset for high-resolution scene-understanding segmentation. In the context of semantic segmentation, this dataset has frequently been used for robustness evaluation, as it addresses real-world and complex outdoor scenarios [rossolini_tnnls_2023, nakka_indirect_2020, rossolini_iccps24]. The Cityscapes validation set includes 500 high-resolution images with an original resolution of ($1024\times 2048$), resized to manage computational costs (as done in [nakka_indirect_2020]). We used this set to evaluate both localized natural corruptions and adversarial attacks.
Models. We targeted multiple models for semantic segmentation, all known for their ability to achieve high performance on large images while maintaining practical inference times affordable for many real-time applications. Specifically, we considered ICNet [icnet_paper], BiseNet [bisenet_paper], DDRNet [ddrnet_paper], PSPNet [pspnet_paper], SegFormer [xie2021segformer], PIDNet [xu2023pidnet], and DeepLabV3 [chen2017rethinking_deeplab], including multiple versions of each model to evaluate robustness differences, such as variations in the backbone architecture. Pretrained versions, available in the paper’s repository, were used for the experimental analysis. The average accuracy of these models on the clean validation set of Cityscapes is reported in the first column of Table 1. For fairness, all results reported for evaluating natural robustness are averaged over 5 runs using the same set of seeds across model comparisons.
Metrics. The spatial robustness analysis was conducted by evaluating the pixel-wise accuracy of the models in the non-corrupted region $A_{\overline{M}}$ and the corrupted region $A_{M}$, both defined in Section 3. Furthermore, to better understand the robustness degradation as the evaluation parameters vary, we analyzed the previous scores in terms of the relative corruption error (RCE) with respect to the accuracy of the addressed models in the same area of interest when no natural corruption or attacks are applied. Leveraging the localized corruption accuracy defined in Equation 4, the relative corruption error can be computed as
$\text{RCE}(f,C,M)=\frac{A_{\mathcal{P}}(f,\{\emptyset\},M)-A_{\mathcal{P}}(f,C,M)}{A_{\mathcal{P}}(f,\{\emptyset\},M)},$ (11)
where $\mathcal{P}_{i}(M)=(1-M_{i})/\overline{\mathcal{S}}$. A similar definition can be derived for the RCE within the corrupted region, considering $\mathcal{P}_{i}(M)=M_{i}/{\mathcal{S}}$. Note also that, as indicated in Section 3, natural corruptions have been evaluated across different regions (see Equation 6).
6.2 Evaluation of Natural Robustness
Type Brightness Snow Gaussian Noise
Severity Clean 1 3 5 1 3 5 1 3 5
Metric MAcc $A_{M}$ $A_{\overline{M}}$ $A_{M}$ $A_{\overline{M}}$ $A_{M}$ $A_{\overline{M}}$ $A_{M}$ $A_{\overline{M}}$ $A_{M}$ $A_{\overline{M}}$ $A_{M}$ $A_{\overline{M}}$ $A_{M}$ $A_{\overline{M}}$ $A_{M}$ $A_{\overline{M}}$ $A_{M}$ $A_{\overline{M}}$
BisenetX39 70.6 63.1 67.0 47.0 61.4 34.7 54.9 34.9 57.4 22.1 51.4 20.6 51.3 50.0 61.5 20.4 50.9 12.1 43.4
BisenetR18 78.0 72.6 76.0 59.5 71.5 49.1 67.9 42.2 68.7 24.7 64.2 19.7 62.0 66.4 73.9 42.4 67.6 21.5 62.4
BisenetR101 82.3 76.8 80.3 63.7 77.3 53.7 74.1 39.7 72.5 20.8 68.2 19.8 68.2 64.0 78.5 35.8 74.3 22.4 70.7
DDRnet23 83.2 78.7 81.7 64.7 76.4 51.2 72.5 39.2 71.8 22.5 67.4 18.5 65.1 64.2 78.1 27.3 67.9 13.7 59.8
DDRnet23Slim 80.8 75.0 78.9 61.4 73.0 49.4 68.8 39.9 68.2 25.4 63.5 20.1 60.4 63.5 74.1 32.9 64.4 15.0 58.3
Icnet 72.7 65.8 70.4 50.3 64.0 41.4 58.6 40.9 58.7 22.7 48.1 20.4 47.6 59.2 67.3 37.5 58.1 20.8 47.8
PSPnet 81.6 75.1 79.1 60.5 74.5 49.1 71.6 35.2 70.1 14.6 63.6 11.6 60.4 66.8 78.2 35.3 70.9 15.6 59.8
Segformer-bo 79.3 75.8 78.2 67.6 76.4 62.0 75.7 45.9 73.2 26.5 70.9 19.6 69.9 68.8 77.1 54.0 74.8 34.8 73.0
Segformer-b1 80.9 77.4 80.2 70.2 79.1 66.0 78.6 50.1 77.3 27.9 75.4 21.1 74.5 72.5 80.2 60.2 79.1 42.2 77.7
Deeplabv3-m 76.7 73.3 75.0 66.6 72.3 59.2 69.4 50.5 69.8 32.1 63.9 25.2 59.9 56.4 71.4 36.3 64.4 20.2 57.6
Deeplabv3-r 80.6 77.0 78.6 71.7 75.7 64.8 73.0 49.3 72.6 29.2 68.4 22.3 66.1 65.8 76.4 40.3 67.6 16.3 58.1
Pidnet-s 80.0 74.1 77.1 61.7 73.4 48.8 69.0 41.7 69.5 22.3 63.9 18.2 61.2 65.6 76.3 34.3 66.2 12.2 52.8
Pidnet-m 80.6 75.9 79.2 62.9 75.2 48.9 71.2 47.6 73.5 28.1 69.6 22.6 67.1 59.4 75.2 27.3 68.8 10.1 64.4
Pidnet-l 82.1 77.2 81.1 60.9 76.7 47.3 72.4 47.7 74.3 24.9 69.3 19.0 67.4 69.0 79.5 36.4 71.7 15.6 63.3
Table 1: Results of the analysis for natural corruptions considering brightness, snow, and gaussian noise as corruption types at different severity levels (1, 2, 3). The first column reports the clean average accuracy (MAcc), while the remaining columns analyze the accuracy within and outside the corrupted region, represented as $A_{M}$ and $A_{\overline{M}}$, respectively, following the analysis presented in Section 3. The corruption ratio is fixed at 50% of the image. For each metric and analysis, the top-2 models are highlighted in green for easier interpretation.
Figure 3: Illustrations of the effects of localized natural corruptions across different semantic segmentation models (BiseNetX39, PSPNet, SegFormer-B1, DeepLabV3-ResNet) with a perturbation ratio of $r=0.3$ and severity level 3. The first row shows the clean image along with its ground truth label and corresponding model outputs. The second row presents the same image after applying localized Gaussian noise, and the third row shows it with synthetic snow applied.
We first evaluate the natural robustness of the selected SS models. Table 1 reports the results for the Non-Corrupted Region Analysis ($A_{\overline{M}}$ in the table) and the Corrupted Region Analysis ($A_{M}$ in the table), considering different localized natural corruptions (synthetic snow, brightness corruptions, and gaussian noise) across different severity levels. For these tests, the corruption ratio is fixed to $r=0.5$, and corruptions are applied by splitting the images into patches of size $(256,256)$.
Considering the robustness in the corrupted region (columns labeled with $A_{M}$), for snow and brightness adjustments, no model demonstrates a clear and consistent superiority over the others, suggesting that convolution-based architectures remain valid alternatives to SegFormers. This observation does not hold in the case of gaussian noise, where transformer-based architectures exhibit a significant accuracy advantage in the corrupted region.
Regarding the robustness in the non-corrupted region (columns labeled with $A_{\overline{M}}$), transformer-like architectures achieve significantly better performance, especially when addressing higher severity levels of localized corruptions. In these cases, natural corruptions have the potential to substantially affect non-corrupted regions, thereby reducing the accuracy of $A_{\overline{M}}$. However, SegFormer demonstrates robustness by preventing the spread of mispredictions to otherwise clean regions.
In Figure 3, we also show the outputs of the models when localized natural perturbations are applied, while the clean image and corresponding outputs of different models are displayed in the first row. For this test, we use a corruption ratio of $r=0.3$ and a severity level of $3$. The corresponding results align with those in Table 1 and previous observations, showing that, in general, the SegFormer architectures achieve better robustness both within and outside the corrupted regions.
(a) Synthetic snow analysis
(b) Gaussian noise analysis
Figure 4: Analysis of the impact of the corruption ratio. We report the relative corruption error RCE with respect to the corrupted region (left plots) and the non-corrupted region (right plots). We used synthetic snow and gaussian noise, in the top and bottom plots, respectively, with severity level 3.
Impact of the corruption ratio
To further analyze the impact of localized corruption, we report in Figure 4(a) and Figure 4(b) the variation of the RCE (computed with respect to the clean accuracy in the addressed region) as a function of the corruption ratio. The RCE was computed both within (left plots) and outside (right plots) the corrupted regions, considering synthetic snow and gaussian noise, respectively.
As shown in the plots, the corruption error follows different trends across the models. For small corruption ratios and corruption types with limited impact, such as synthetic snow, convolution-based models such as DeepLab-MobileNet and DeepLab-ResNet achieve performance comparable to SegFormer, both within and outside the corrupted region. However, when the corruption ratio increases, the error rises more rapidly for the convolution-based models, while SegFormer demonstrates superior robustness. As discussed above, a different behavior is observed when considering gaussian noise, for which SegFormer already exhibits significantly higher robustness even for small corrupted regions. We believe that this advantage is largely attributed to the application of global attention in transformers, which can effectively attend to uncorrupted areas by leveraging features from those regions or, when regions are completely perturbed, distribute robust features more efficiently across the image [gu2022vision].
(a) Attack with patches (100,100)
(b) Attack with patches (200,200)
Figure 5: Analysis of the accuracy in the non-corrupted region with (dark color) and without (light colors) localized adversarial perturbations. Attacks were performed using $\epsilon=16/255$ and $\epsilon=32/255$. The patch is always applied at the center of the image, with a size of (100, 100) in (a) and (200, 200) in (b), respectively.
(a) Segformer Attack
(b) PSPNet Attack
Figure 7: Illustration of the impact of using the Region-aware multi-attack adversarial analysis (Algorithm 1) on Segformer (a) and PSPNet (b) for two different images in the Cityscapes validation set. The use of multiple attacks allows for coverage of different parts of the input pixels, highlighting the complexity of solving a multi-objective adversarial optimization problem for SS and demonstrating how the proposed algorithm can overcome this challenge. The attacks shown here use patches of size (200, 200) with $\epsilon=32/255$.
6.3 Evaluation of Adversarial Robustness
To further understand localized robustness, we evaluate the performance of the SS models using the proposed region-aware multi-attack adversarial analysis (Algorithm 1), aiming for the identification of the worst-case scenario for spatial robustness.
As discussed in Section 5, for adversarial evaluation, it is important to note that we focus on the analysis of non-corrupted regions ($A_{\overline{M}}$). This is because evaluating corrupted regions under localized attacks is straightforward, as any model will typically exhibit near-zero accuracy within the corrupted area. In contrast, analyzing the non-corrupted regions provides a more meaningful assessment of spatial robustness.
Please note that, for completeness, concerning the adversarial analysis, the full list of parameters involved is: the patch size, the attack magnitude $\epsilon$, the number of attacks $N_{\text{Att}}$, the position of the attack (which is always centered unless explicitly stated otherwise), and the number of iterations for each attack.
In Figure 5(a) and Figure 5(b), we report the accuracy in the non-corrupted regions under attack, compared to the non-attacked case (‘clean’, shown in the lightest color in the bars), using localized adversarial attacks with patch sizes of (100, 100) and (200, 200), centered in the image, respectively. In both cases, we tested attacks with $\epsilon$ values of $16/255$ and $32/255$ (indicated by different colors in the plots). The multi-attack of Algorithm 1 is performed with $N_{\text{Att}}=3$ attacks and 50 optimization iterations (see Equation 10) per attack. Further analysis of these parameters is discussed in the following.
As shown in the plots, the trend is completely opposite to what we observed for natural corruptions. In this case, transformers exhibit lower robustness, primarily due to the global attention mechanisms they employ [gu2022vision, xie2021segformer, vaswani_attention_2017]. For example, even with a small patch size (100, 100), the accuracy drops significantly from 0.73 to 0.25 and 0.17 for $\epsilon=16/255$ and $\epsilon=32/255$, respectively. In contrast, convolution-based models, such as DeepLabV3 with ResNet backbone, demonstrate significantly higher robustness. This robustness can be attributed to the limited local receptive field inherent to convolutional layers [luo2016understanding]. However, it is important to note that these models are not purely convolutional, as they also integrate attention mechanisms to enhance performance. This integration can result in substantial accuracy drops, as seen with PSPNet when using patches of size (200, 200). This observation is consistent with previous studies, which suggest that even convolutional models employing high-context exploitation through attention-based mechanisms (such as pyramidal squeezing and excitation blocks) can still be vulnerable to certain localized adversarial perturbations [rossolini_tnnls_2023, nakka_indirect_2020].
Figure 8: Analysis of the impact of the proposed localized adversarial framework under different attack numbers $N_{\textit{Att}}$ in Algorithm 1 (a), and under different number of iterations (b). In (a), the number of iterations is fixed to 50, while in (b), the number of attacks is fixed to 3. In both analyses, patches are of fixed size (200, 200) and $\epsilon=16/255$.
Benefits of Multi-Attacks
To better highlight the benefits provided by Algorithm 1, we present in Figure 8(a) the increase of the RCE, measured with respect to the clean accuracy in the non-corrupted regions, as the number of attacks $N_{\text{Att}}$ in Algorithm 1 increases. For these tests, we used patches of size (200, 200), $\epsilon=16/255$, and 50 optimization iterations per attack.
Specifically, with a single attack (#1), the multi-attack is equivalent to running a standard attack over the entire targeted region. However, as the number of attacks increases, the algorithm begins focusing on fooling regions that have not yet been attacked, thereby expanding the coverage to new regions that were previously unaffected. In fact, as one may expect, the effectiveness of the multi-attack significantly improves with more attacks, as indicated by the increase in RCE. This effect is particularly pronounced in models with high attention mechanisms, such as transformers and PSPNet, demonstrating the importance of the multi-attack approach in identifying worst-case scenarios of localized adversarial perturbations. For example, the RCE of SegFormer models increases significantly, rising from approximately 0.6 with a single attack to around 0.95 when using five attacks ($N_{\text{Att}}=5$).
We also provide illustrations of the effect of multiple attacks for SegFormer and PSPNet in Figures 7(a) and 7(b). As shown in the plots, the adoption of multiple attacks with awareness of the fooling regions allows for overcoming issues in the multi-objective adversarial problem for SS. In Figure 7(a), at each attack, the untargeted multi-attacks target different classes, highlighting the inter-class problem discussed previously in Figure 2(a). Similarly, for PSPNet (Figure 7(b)), addressing different fooling regions leads to a third attack capable of affecting the entire image, thereby overcoming potential objectives that previously limited the effectiveness of earlier attacks.
Impact of the Number of Optimization Iterations
We also analyze in Figure 8(b) the impact of varying the number of optimization iterations used at line 9 in Algorithm 1, fixing the number of attacks $N_{\text{Att}}=3$. As expected, the RCE increases when using more optimization iterations, such as 200 or 250, highlighting that the complexity of the multi-objective optimization problem in this context requires a higher number of iterations to improve effectiveness. However, we acknowledge that for large-resolution datasets, as Cityscapes, a limited number of iterations (e.g., 50) is sufficient to provide meaningful and consistent comparisons, as demonstrated in the plots. This balance helps find a trade-off between computational cost and robustness evaluation, particularly when dealing with high-resolution images, which are computationally and memory expensive due to the gradient computations required to update the attack at each step.
Figure 9: Differences in relative corruption error when applying localized adversarial attacks at the center of the image versus the bottom-left corner. Patches of size (200, 200) and $\epsilon=16/255$ are considered.
On the Position of the Attacks
Finally, we highlight the importance of the position of the adversarial perturbed patch within the image. The previous attacks were conducted with the patch placed at the center of the image, which represents the worst-case position for standard convolutional models due to its ability to affect the largest possible receptive field [luo2016understanding, rossolini_iccps24, nakka_indirect_2020]. However, it is important to analyze how the effectiveness of localized adversarial attacks decreases when the patch is shifted to more constrained areas, such as the corners of the image.
In Figure 9, we compare the difference between the RCE when attacking the central area versus the one when attacking the bottom-left corner. Patches of size (200, 200) and $\epsilon=16/255$ were considered. Specifically, the larger the difference, the more the model demonstrates robustness with respect to the attack position.
As expected, transformer-based architectures, which rely on global attention and process images through patches, exhibit lower sensitivity to the position of adversarial patches. Even when the patch is moved away from the center, the model’s response remains close to the worst-case scenario. This highlights the critical risk of attacks for such architectures, being effective from multiple locations within the image, leading to the same adversarial outcome. In contrast, convolutional models exhibit more variable behavior depending on the patch’s position. However, since the tested models are not purely convolutional and incorporate attention mechanisms, their robustness is still partially affected when the patch is placed in a corner. For example, BiSeNet with a ResNet-101 backbone and PSPNet demonstrate differences of approximately 0.05 and 0.06 in RCE between central and corner positions, respectively, indicating that localized attacks can still significantly degrade their performance.
(a) $f_{1}$: DeepLab_MobileNet
$f_{2}$: SegFormer._b0
(b) $f_{1}$: DeepLab_MobileNet
$f_{2}$: SegFormer._b1
(c) $f_{1}$: DeepLab_ResNet
$f_{2}$: SegFormer._b0
(d) $f_{1}$: DeepLab_ResNet
$f_{2}$: SegFormer._b1
(e) $f_{1}$: Bisenet-R101
$f_{2}$: SegFormer._b0
(f) $f_{1}$: Bisenet-R101
$f_{2}$: SegFormer._b1
(g) $f_{1}$: DDRNet23
$f_{2}$: SegFormer._b0
(h) $f_{1}$: DDRNet23
$f_{2}$: SegFormer._b1
Figure 11: Analysis of the natural and adversarial errors, computed as discussed in Equations 13 and 14, respectively, and clean accuracy when considering different models as $f_{1}$ and $f_{2}$ in the ensemble strategy defined in Equation 12, for different configurations of $\gamma$. The results aim to explore the existence of a possible trade-off to balance robustness against localized natural corruptions and adversarial perturbations.
6.4 Searching for a Trade-off in Ensembling
Given the results obtained in the above analyses of natural perturbations and adversarial attacks, we acknowledge that the selected transformer architectures generally exhibit greater robustness to localized natural corruptions by effectively reducing misclassifications, not only within the corrupted regions but also in the surrounding non-corrupted areas. However, their reliance on high attention mechanisms makes them more vulnerable to localized adversarial attacks, resulting in a significant drop in accuracy compared to convolutional models tested.
Following this observation, we emphasize the importance of searching for a trade-off that balances natural robustness and worst-case adversarial robustness, which could be particularly beneficial in safety-critical domains such as autonomous driving. To address this challenge, we investigate the use of model ensembling, with natural and adversarial localized robustness as key metrics to evaluate the effectiveness of the ensemble strategy. To explore this approach, we first define a test-time weighted ensemble strategy [ensemble_GANAIE2022105151] as:
$g_{\gamma}(f_{1},f_{2})=\gamma\cdot f_{1}+(1-\gamma)\cdot f_{2},$ (12)
where $f_{1}$ and $f_{2}$ are two models selected for the ensemble, and $\gamma$ is a parameter that balances the importance of the softmax scores of model $f_{1}$ with respect to $f_{2}$ in the final model $g$.
Specifically, for our tests, we analyze pairs of models where $f_{1}$ is a model with higher robustness to localized adversarial attacks (e.g., DeepLab-ResNet101, DeepLab-MobileNet, BiSeNet-ResNet101, and DDRNet23), and $f_{2}$ demonstrates stronger robustness to natural corruptions but lower robustness to localized adversarial attacks (e.g., SegFormer_b0 or SegFormer_b1).
To understand the benefits of ensembling, we redefine the RCE for natural and adversarial perturbations, as follows, to account for the combination of $f_{1}$ and $f_{2}$. In particular, most important to us for this analysis, we used as metrics: (i) the RCE (Equation 11) for natural perturbations within the corrupted area, i.e.,
$\textit{CE}_{\textit{Nat}}(g,f_{2},C,M)=\frac{A_{\mathcal{P}}(f_{2},C,M)-A_{\mathcal{P}}(g,C,M)}{A_{\mathcal{P}}(f_{2},C,M)},$ (13)
where $\mathcal{P}_{i}(M)=M_{i}/\mathcal{S}$ and $C$ is a set of natural corruptions (we considered gaussian noise, with $r=0.5$); and (ii) the RCE for adversarial perturbations outside the corrupted area, i.e.,
$\quad\textit{CE}_{\textit{Adv}}(g,f_{1},C,M)=\frac{A_{\mathcal{P}}(f_{1},C,M)-A_{\mathcal{P}}(g,C,M)}{A_{\mathcal{P}}(f_{1},C,M)},$ (14)
where $\mathcal{P}_{i}(M)=(1-M_{i})/\overline{\mathcal{S}}$ and $C$ is a localized adversarial corruption (we used the attack proposed in Algorithm 1, with $N_{\textit{Att}}=5,\epsilon=16/255$).
The rationale behind these error definitions is that, since $f_{1}$ and $f_{2}$ are chosen as the best-performing models for natural and adversarial robustness, respectively, this formulation allows us to understand how varying $\gamma$ influences improvements in natural robustness and the corresponding drop in adversarial robustness.
The results of Figure 11 confirm the expected trend. As we increase $\gamma$, the ensemble model becomes more robust to natural perturbations but also increasingly susceptible to localized adversarial attacks. Note that, in this case, the attacks are performed directly on the ensemble model to provide a fair evaluation, considering the corresponding $\gamma$ value. Interestingly, for certain values (e.g., $\gamma=0.4$), the adversarial robustness is largely preserved, while we observe a significant reduction in the natural error. This phenomenon is particularly noticeable when DeepLab architectures are used as $f_{1}$, as the natural error decreases slightly before the adversarial error curve begins to rise. This demonstrates that a positive trade-off between natural and adversarial robustness can be found.
Additionally, we report the mean accuracy of the ensemble on clean inputs in the top portions of the plots to ensure fairness and confirm that there is no drop in clean accuracy. The error of clean accuracy is computed with respect to the best-performing model between $f_{1}$ and $f_{2}$. As shown in the plots, the use of a well-selected $\gamma$ gets also to negative errors, indicating that the ensemble strategy improves performance on the clean validation set.
We believe this analysis highlights the importance of rethinking model ensembling to achieve results beyond raw performance. This approach can be particularly effective when combining architectures with orthogonal strengths, such as robustness to localized natural corruptions and robustness to localized adversarial attacks.
7 Discussion and Conclusion
This work presented an extensive analysis of the robustness of segmentation models under localized corruptions, considering both natural and adversarial corruptions. We first introduced metrics that extend the classic pixel-wise accuracy, commonly used in segmentation, to provide a more comprehensive analysis of spatial robustness, both within and beyond the corrupted regions. We then proposed a framework that enables a systematic evaluation of localized natural and adversarial corruptions by allowing the selection of parameters, including corruption ratios and model patch sizes.
Our analysis of adversarial perturbations also improves upon existing strategies in the literature by demonstrating that, in dense prediction tasks such as semantic segmentation, targeting multiple pixels introduces a complex multi-objective problem. This complexity makes it difficult to assess robustness in the worst-case adversarial scenario using a single localized perturbation. Therefore, we proposed the region-aware multi-attack adversarial analysis, a method to assess the spatial robustness of models by applying multiple $\infty$-norm attacks while continuously adjusting the fooling area.
All findings were evaluated across a diverse set of segmentation models in a driving scenario, specifically using the Cityscapes dataset [cordts2016cityscapes]. Our results revealed distinct differences in model behavior under natural and adversarial corruptions. Furthermore, we highlighted the importance of finding a trade-off between these two robustness aspects, which are not necessarily correlated. As a preliminary step toward addressing this challenge, we explored ensemble strategies that combine convolution-based models with transformer-based architectures.
In a future work, we aim at further investigating this trade-off by exploring more complex ensembling approaches, extending the analysis to other tasks and types or forms of corruptions, and providing theoretical insights into this intriguing mixed property. Another important direction is to integrate the proposed localized robustness analysis into model training strategies through the use of localized augmentation techniques (e.g., Cutout [devries2017improved_cutout] and CutMix [yun2019cutmix]), which are rarely applied in complex scene understanding tasks such as semantic segmentation. This could provide valuable insights into the development of novel defense mechanisms and robust training approaches. In this work, we focused on standalone models trained with conventional image-level augmentations, leaving this integration as a promising direction for future research.
To conclude, our study represents an important advancement in understanding the spatial robustness of segmentation models, while also identifying key challenges and open research directions for future investigations.
Acknowledgements
This work was partially supported by project SERICS (PE00000014) under the MUR (Ministero dell’Università e della Ricerca) National Recovery and Resilience Plan funded by the European Union - NextGenerationEU.